### v1

Abstract
Despite the popularity of large language model (LLM) quantization for inference acceleration, significant uncertainty remains regarding the accuracy-performance trade-offs associated with various quantization formats. We present a comprehensive empirical study of quantized accuracy, evaluating popular quantization formats (FP8, INT8, INT4) across academic benchmarks and real-world tasks, on the entire Llama-3.1 model family. Additionally, our study examines the difference in text generated by quantized models versus their uncompressed counterparts. Beyond benchmarks, we also present a couple of quantization improvements which allowed us to obtain state-of-the-art accuracy recovery results. Our investigation, encompassing over 500,000 individual evaluations, yields several key findings: (1) FP8 weight and activation quantization (W8A8-FP) is lossless across all model scales, (2) INT8 weight and activation quantization (W8A8-INT), when properly tuned, incurs surprisingly low 1-3% accuracy degradation, and (3) INT4 weight-only quantization (W4A16-INT) is competitive with 8-bit integer weight and activation quantization. To address the question of the ”best” format for a given deployment environment, we conduct inference performance analysis using the popular open-source vLLM framework on various GPU architectures. We find that W4A16 offers the best cost-efficiency for synchronous deployments, and for asynchronous deployment on mid-tier GPUs. At the same time, W8A8 formats excel in asynchronous “continuous batching” deployment of mid- and large-size models on high-end GPUs. Our results provide a set of practical guidelines for deploying quantized LLMs across scales and performance requirements.
footnotetext: Correspondence to: $<$firstname$>$@neuralmagic.com.
1 Introduction
The massive computational costs of serving large language models have led to a significant amount of work on inference acceleration techniques, such as quantization (Frantar et al., 2022; Dettmers & Zettlemoyer, 2022; Lin et al., 2024a), speculative decoding (Chen et al., 2023; Leviathan et al., 2023), or model pruning (Xia et al., 2023; Muralidharan et al., 2024). Model quantization, i.e., reducing the bitwidth of weights, activations, or both, is arguably the standard approach to reducing memory and computational costs at deployment time. The critical trade-off in LLM quantization is between the speed/memory benefits and the accuracy of the resulting compressed model. Despite extensive research on quantization, few studies provide systematic benchmarks or practical guidelines on expected performance at different compression levels. This state of affairs is expressed in our fore-title, which uses a direct quote from the wave of user feedback regarding the quantized Llama-3.1-405B model (Dubey et al., 2024), initially suspected to drop significant accuracy (Zhang, 2024; Karpathy, 2024) and later shown to be near-lossless in LMSYS Arena user testing (Chiang et al., 2024). This uncertainty can be a significant barrier to adoption of quantized formats. In this paper, we systematically consider the following question:
What are the practical accuracy-performance trade-offs for popular quantization formats?
Due to the broad range of existing quantization techniques, we focus our study on formats with efficient computational support across different batching levels and low accuracy loss. This reduces the range of formats to be examined to 8-bit weights and activations (W8A8), in integer (INT) precision for NVIDIA Ampere and older devices, and floating-point (FP) precision for NVIDIA Hopper and Ada Lovelace, as well as 4-bit integer weights with 16-bit activations (W4A16-INT). These formats are efficiently supported in the vLLM inference engine (Kwon et al., 2023), which we use for performance experiments through the GuideLLM framework Neural Magic (2024).
To examine accuracy, we implemented a broad, automated evaluation suite, including leading academic benchmarks, namely the Open LLM Leaderboard V1 (Beeching et al., 2023), and its recent, more challenging, V2 version (Fourrier et al., 2024). We complement this examination with a set of real-world benchmarks, which focus on generative quality in an open-ended setup, specifically Arena-Hard-Auto-v0.1 Li et al. (2024b), HumanEval Chen et al. (2021), and HumanEval+ Liu et al. (2023a). We further extend these benchmarks with an in-depth text similarity analysis between outputs from uncompressed and quantized models. Finally, we examine the performance implications of deploying quantized LLMs. For this, we benchmark across all quantization formats in vLLM (version 0.6.1.post2) across three GPU types (A6000, A100, H100) in six latency-sensitive use cases and two throughput use cases. Our main findings are as follows:
  1. 1.
    W8A8-FP quantization is essentially lossless: Across all benchmarks, it preserves the uncompressed model’s average accuracy or falls within the evaluation’s margin of error. Remarkably, we achieve this via a simple and robust approach: dynamic per-token activation quantization and symmetric weight quantization via round-to-nearest assignment.
  2. 2.
    W8A8-INT quantization shows a tolerable accuracy loss, in the 1-3% per-task average range. This is achieved using dynamic activation quantization or the more complex SmoothQuant technique (Xiao et al., 2022); for weights, we use GPTQ (Frantar et al., 2022) with symmetric quantization. Our findings stand in contrast with prior work (Li et al., 2024a; Lee et al., 2024b), which reported large accuracy drops (more than 10% on average) in the same setting.
  3. 3.
    W4A16-INT quantization maintains consistently low accuracy loss, being competitive in terms of accuracy to W8A8-INT. Our findings are obtained for a simple variant of GPTQ, which we surprisingly find to be more accurate on real-world tasks than the more recent AWQ method Lin et al. (2024a).
  4. 4.
    On the text analysis side, we show that, in autoregressive text generation, larger quantized models tend to closely follow the original word choices and sentence structures of the baseline. At the same time, on smaller models, quantization introduces moderate structure variability relative to full precision, but still preserves semantic meaning.
  5. 5.
    For quantized inference performance, we show that W4A16-INT offers favorable performance in synchronous deployments. For asynchronous deployments, W4A16 is preferable for small models on mid-tier GPUs, and W8A8 formats are more effective for mid-to-large-size models on high-end GPUs. The most cost-effective quantization scheme depends on model size, hardware, use case (e.g., code completion vs. multi-turn chat), and deployment scenario (synchronous vs. asynchronous).
Overall, this work provides a first in-depth study of accuracy-vs-performance-vs-cost trade-offs for quantized models across various formats, algorithms, use cases, and hardware types. We hope it can serve as a guide for practitioners looking to deploy models efficiently, and as motivation to researchers interested in method improvements.
2 Background and Related Work
2.1 A Primer on Quantization
In the following, we provide some background on post-training quantization of LLM weights and activations. Early work on this topic has considered INT8 activation quantization and INT4/INT8 weight quantization (Dettmers et al., 2022; Yao et al., 2022; Park et al., 2022). A standard approach for weight quantization has been direct round-to-nearest (RTN) quantization over small groups: given a group of $g$ consecutive weights, viewed as a vector $\mathbf{x}\in\mathbb{R}^{g}$, $b$-bit RTN is defined as:
${\cal Q}(\mathbf{x},b)$ $=\operatorname{rnd}\Bigg{(}\frac{\mathbf{x}-\min(\mathbf{x})}{\max(\mathbf{x})-\min(\mathbf{x})}(2^{b}-1)\Bigg{)}$
$=\operatorname{rnd}((\mathbf{x}-z(\mathbf{x}))/s(\mathbf{x})),$ (1)
where $\operatorname{rnd}$ rounds to the nearest integer, $z=z(\mathbf{x})=\min(\mathbf{x})$ is the “zero point” and $s=s(\mathbf{x})=(\max(\mathbf{x})-\min(\mathbf{x}))/(2^{b}-1)$ is the scale, taken as min-max. This early work raised two key issues: RTN on weights loses significant accuracy for INT4 precision, and RTN quantization of activations even to higher INT8 precision can be lossy due to the large “outlier features” in LLMs Dettmers et al. (2022).
Weight quantization. Towards the first issue, GPTQ (Frantar et al., 2022) improved upon RTN for scalar, per-weight, quantization by allowing weights to be adjusted via second-order updates computed over a sample of calibration data. Follow-up scalar quantization methods such as AWQ (Lin et al., 2024a), SqueezeLLM (Kim et al., 2023), OWQ (Lee et al., 2024a), and SpQR (Dettmers et al., 2023) implemented variants of outlier-aware quantization, where a small fraction of weights are effectively stored in higher precision. These methods can reach highly-accurate 4-bit quantization, but the latter three lack general kernel support.
More recently, several high-compression methods have been proposed, such as QuIP (Chee et al., 2023), QuIP# Tseng et al. (2024a), QTIP (Tseng et al., 2024b) AQLM (Egiazarian et al., 2024), and GPTVQ (van Baalen et al., 2024), targeting the 3-bit range or lower. These methods leverage much more complex representations, such as lattice/vector quantization, often paired with incoherence pre-processing of the weights. Despite high accuracy, these methods add overheads and lack efficient kernel support for batch sizes larger than 1, which is necessary for general deployments. Thus, our study does not currently consider these formats.
Activation quantization. Quantizing both activations and weights to 8-bit or 4-bit precisions allows direct utilization of existing hardware support for lower-bit operations. However, activation matrices are difficult to quantize effectively (Dettmers et al., 2022) due to the presence of outlier features — elements that can be up to 100 times larger than the average. Dettmers et al. (2022) extracted columns at runtime and quantized the rest to INT8, which is inefficient. SmoothQuant (Xiao et al., 2022) improves upon this by recognizing that outlier features are generally stable across the model and can be extracted utilizing a calibration set and folded during compilation.
Follow-ups to SmoothQuant investigate the possibility of W4A4 quantization (Ashkboos et al., 2023; 2024) or mixed-precision W4A8 (Lin et al., 2024b; Zhang et al., 2024), also considering quantizing the KV-cache. While promising, these approaches still lose significant accuracy and lack efficient support in high-performance inference frameworks. We, therefore, also leave their investigation for future work.
2.2 Efficient Inference With Quantized LLMs
Several performant LLM inference frameworks exist, such as HuggingFace TGI (HuggingFace, 2024), NVIDIA TensorRT-LLM (NVIDIA, 2023), and vLLM (Kwon et al., 2023). We focus our benchmarking efforts on vLLM, a popular open-source framework that addresses the specific challenges of serving large, autoregressive neural networks.
The backend architecture of vLLM combines efficient execution management with high-performance computation graphs that utilize pre-compilation, optimized operator fusion, and performant scheduling. Additionally, vLLM incorporates computational acceleration through customized kernels leveraging NVIDIA CUTLASS and hand-written GPU kernels. At the same time, vLLM enables broad hardware compatibility, supporting NVIDIA and AMD GPUs, TPUs, and CPUs. Several optimization techniques, including FlashAttention (Dao et al., 2022), FlashInfer (FlashInfer, 2023), PagedAttention (Kwon et al., 2023), and Marlin (Frantar et al., 2024) are also supported. This support and its open-source nature have led to vLLM’s adoption by industry and the open-source community. The framework integrates optimized CUTLASS kernels for W8A8 (FP and INT), and mixed-precision Marlin kernels for W4A16-INT.
2.3 Accuracy Evaluations of Quantized LLMs
There is a substantial amount of prior work on the evaluation of quantized LLMs, primarily focusing on accuracy trade-offs under various quantization schemes Yao et al. (2023); Liu et al. (2023b); Huang et al. (2024); Gong et al. (2024b); Li et al. (2024a); Gong et al. (2024a). However, much of this research has concentrated solely on academic benchmarks, which do not fully represent the real-world environments where LLMs are deployed. Furthermore, the fact that some of these works do not tune hyperparameters and algorithmic choices for the quantization algorithms can lead to incorrect conclusions regarding accuracy, as we demonstrate later in the experimental section. Specifically, we refer to the claim that 8-bit integer activation quantization causes significant accuracy drops Li et al. (2024a); Lee et al. (2024b), which we contradict experimentally.
The closest work to ours is by Lee et al. (2024b), which, like most prior studies, focuses solely on accuracy under quantization. However, we identify several points of divergence. First, although the authors claim to analyze models up to 405B in size, open-ended benchmarks are not included for this scale, and accuracy scores for the full-precision uncompressed model are missing even for academic benchmarks. Without baselines, it is difficult to draw meaningful conclusions on the impact of quantization at this scale. To address this, we enabled highly efficient multi-node evaluations for the 405B model, allowing comprehensive benchmarking in full precision across academic and real-world tasks.
The second divergence point concerns their claims that AWQ outperforms GPTQ in a 4-bit weight-only quantization setup. We attribute this finding to suboptimal hyperparameter choices in GPTQ quantization and provide a detailed comparison of the methods in Table 1 and Appendix A.2. Our results show that both methods perform similarly on academic benchmarks, while GPTQ demonstrates notable improvements over AWQ on real-world benchmarks, especially in coding tasks.
Third, we also challenge the prior conclusion that W8A8-INT performs substantially worse than W8A8-FP and W4A16-INT. When tuned carefully, W8A8-INT proves competitive, with only minor accuracy losses. For instance, while Lee et al. (2024b) reports a 10-point drop in accuracy for the W8A8-INT quantized 405B model on the Open LLM Leaderboard V2 benchmark relative to FP8 quantized one, our carefully optimized quantization parameters reduce this gap to just 0.7 points.
Unlike previous work, we also explore how quantization affects text generation quality. This analysis provides deeper insights into the impacts of various quantization schemes on sentence structure, word choice, and semantic fidelity in text generated by quantized LLMs.
Furthermore, we conclude our study by analyzing the performance impacts of deploying various quantization formats across a range of scenarios that represent real-world applications. This examination provides insights into how each quantization approach performs under practical deployment conditions and hardware configurations.
3 Benchmark Design and Setup
3.1 Datasets and Benchmarks
Our accuracy evaluation suite is designed to encompass a wide range of inference scenarios and use cases. It includes a diverse set of tasks to ensure a representative collection of topics relevant to practical deployments. Furthermore, the evaluation process is designed to be completely automated, as opposed to human-in-the-loop approaches, enabling streamlined experimentation. Overall, we group all of the benchmarks into three distinct categories.
Academic benchmarks such as Open LLM Leaderboard V1 and V2 Beeching et al. (2023); Fourrier et al. (2024), are essential for evaluating model improvements. They focus on structured tasks like question-answering and reasoning, providing consistent and reproducible scoring. However, they often lack alignment with real-world scenarios where semantics, variability, and context-awareness are important.
The Open LLM Leaderboard V1 consists of a diverse range of topics, including grade school math (GSM Cobbe et al. (2021)), world knowledge and reasoning (MMLU Hendrycks et al. (2020), ARC-Challenge Clark et al. (2018)), language understanding (Winogrande Sakaguchi et al. (2021), HellaSwag Zellers et al. (2019)), and truthfulness (TruthfulQA Lin et al. (2021)).
The Open LLM Leaderboard V2 builds on this with topics such as expert knowledge and complex reasoning (MMLU-Pro Wang et al. (2024), GPQA Rein et al. (2023), Big Bench Hard Suzgun et al. (2022)), multi-step reasoning (MuSR Sprague et al. (2024)), advanced math problems (MATH Level 5 Hendrycks et al. (2021)), and instruction following (IFEval Zhou et al. (2023)).
By evaluating with both Leaderboards, we cover a broad spectrum of topics. We employ log-likelihood (multiple-choice) and text-generation evaluations, allowing for stress testing of quantized LLMs under diverse conditions.
Real-world benchmarks, unlike academic benchmarks, test models in scenarios that mimic human usage, such as instruction following, chat, and code generation. These benchmarks include Arena-Hard-Auto-v0.1 Li et al. (2024b); Chiang et al. (2024); Li et al. (2024c), HumanEval Chen et al. (2021), and HumanEval+ Liu et al. (2023a), which offer a broader range of tasks with higher variation but better reflect real-world environments. The LMSYS Chatbot Arena Chiang et al. (2024) has established itself as a leading benchmark for LLMs, assessing how models align with human preferences. Arena-Hard-Auto-v0.1 is an automated extension of this benchmark, where an LLM judges responses to 500 complex prompts on various topics. It has demonstrated a strong correlation with human evaluations, achieving a state-of-the-art 89% agreement with human preference rankings Li et al. (2024c). This enables a reliable, fast, and automated evaluation of the model’s chat capabilities without requiring a human. In addition to chat-based interactions, LLMs are widely deployed as coding assistants. To evaluate the performance of quantized models in code generation, we tested them on HumanEval and its more challenging variant, HumanEval+. These benchmarks measure a model’s ability to generate correct and functional code based on programming problems, with HumanEval+ introducing more rigorous evaluation protocol.
Text similarity analysis evaluates how closely the generated outputs of quantized models align with those of their unquantized counterparts. While real-world benchmarks reflect practical usage scenarios for LLMs, their inherently open-ended format introduces notable variance in results. To address this, we extend our evaluation suite to include an analysis of output similarity under identical prompt requests. This allows us to get insights into how quantized models’ text generation differs from that of full-precision models. For this investigation, we employ metrics such as ROUGE Lin (2004), BERTScore Zhang et al. (2019), and Semantic Textual Similarity (STS) to assess both semantic and structural consistency, ensuring that the intended meaning and quality of generated text are preserved under quantization.
ROUGE-1 measures word-level (unigram) overlap between outputs from quantized and unquantized models, while ROUGE-L captures structural similarity by focusing on the longest common subsequence. BERTScore evaluates token-level contextual similarity, computed using the RoBERTa-large model’s contextual embeddings. STS assesses overall semantic similarity at the sentence level, using the Sentence Transformers framework Reimers & Gurevych (2019) built on the 6-layer MiniLM architecture Wang et al. (2020).
With this extensive evaluation framework, we ensured that deployment scenarios ranging from structured, research-driven tasks to open-ended, real-world applications were covered, providing a holistic view of the performance and capabilities of quantized LLMs.
3.2 Models, Formats, and Quantization Algorithms
We focus our evaluation on extensive testing of the recent Llama 3.1 series of models Dubey et al. (2024), which have gained significant traction in research and deployment contexts. To obtain a comprehensive understanding of trade-offs, we conduct all of our experiments on the instruction-tuned versions of all three available Llama 3.1 sizes (8B, 70B, and 405B), and for each of these, we focus on investigating three popular quantization formats with efficient inference support in vLLM: W8A8-FP, W8A8-INT, and W4A16-INT.
Table 1: Comparison of GPTQ and AWQ 4-bit weight quantization algorithms (W4A16-INT). We observe a small gap between methods on academic benchmarks (left) but a more pronounced difference in favor of GPTQ on real-world (open-ended) benchmarks (right).
Academic Benchmarks Real-World Benchmarks
Model Average Score Leaderboard V1 Leaderboard V2 Average Score Arena-Hard HumanEval MBPP
Llama-3.1-8B-Instruct 50.84 74.06 27.62 53.7 25.8 67.3 68.1
GPTQ Frantar et al. (2022) 49.82 73.11 26.53 52.3 24.0 67.1 65.8
AWQ Lin et al. (2024a) 50.05 72.69 27.40 49.4 22.3 63.0 62.8
Llama-3.1-70B-Instruct 62.93 84.20 41.66 73.1 57.0 79.7 82.5
GPTQ Frantar et al. (2022) 62.18 83.77 40.58 73.1 57.0 80.5 81.9
AWQ Lin et al. (2024a) 62.53 83.96 41.09 72.3 56.7 79.4 80.8
W8A8-FP This scheme uses an 8-bit floating point format for weights and activations of all linear operators within transformer blocks. We compress models to this format using a simple round-to-nearest algorithm. Weight quantization follows a symmetric per-output-channel scheme, while activations are dynamically quantized per token. Due to the algorithm’s simplicity, this format requires no calibration data and is efficient even for massive models.
W8A8-INT This scheme reduces the weights and activations of all linear operators within transformer blocks to 8-bit integer values. We apply symmetric per-output-channel GPTQ quantization for weight compression, and for activations, we use dynamic per-token quantization. We found that this scheme performs consistently well across 8B and 405B models, though it results in noticeable accuracy drops for the 70B model. To address this, we employ SmoothQuant, which redistributes the complexity of activation quantization onto the generally easier-to-quantize weights. For calibration data, we found that random tokens are adequate for the 8B scale; however, for larger models, we found that using the high-quality OpenPlatypus dataset Lee et al. (2023) is crucial for consistent accuracy recovery of competitive scores on academic benchmarks.
W4A16-INT In this scheme, the weights of all linear operators within transformer blocks are quantized to 4-bit integers, while activations are retained at 16-bit precision. Weights are compressed using GPTQ quantization applied to groups of 128 consecutive elements with MSE-optimal clipping factors. Due to higher compression, using random tokens proved detrimental to accuracy; therefore, we employ calibration data sampled from the OpenPlatypus dataset.
INT4 Quantization Algorithms. As shown by prior work Dettmers et al. (2022); Lin et al. (2024a); Xiao et al. (2022), round-to-nearest performs poorly for integer weight quantization. Therefore, we focus on two very popular techniques with efficient inference support in vLLM: AWQ and GPTQ. We examine their performance on several benchmarks: Leaderboard V1 and V2, Arena-Hard, HumanEval, and MBPP. We report results in Table 1, and provide detailed per-task breakdown of scores in Appendix Tables 1213, and  8. We observe that, on average, on academic benchmarks, the methods are essentially tied. The advantage of AWQ, on a scale of 0–100, is only 0.23 and 0.35, respectively. However, we found that, on real-world benchmarks, GPTQ consistently outperforms AWQ by larger margins of 2.9 and 0.8, respectively. Consequently, we adopt GPTQ as our primary method for weight quantization. This conclusion is slightly at odds with prior work (Lin et al., 2024a; Huang et al., 2024) observing results that are either in favor of AWQ, or tied across a sub-sample of academic benchmarks. This is explained as follows: (1) contrary to this prior work, we use the version of GPTQ with MSE-optimal clipping factors (as opposed to min-max optimal), which has no overhead and performs consistently better; (2) we use different, higher-quality calibration data; (3) we evaluate across real-world benchmarks.
4 Quantization versus Accuracy
We begin our discussion by examining the accuracy of quantized models across Leaderboard V1 (Table 2), Leaderboard V2 (Table 3) and real-world benchmarks (Table 3). Given the density of the results, we examine them individually via average recoveries across higher-level benchmarks and discuss “outlier” observations.
Table 2: Detailed per-task breakdown of accuracy on a subset of academic benchmarks (Open LLM Leaderboard V1) for quantized Llama-3.1-Instruct models across all three model sizes (8B, 70B, 405B). Higher score is better.
Recovery % Average Score MMLU 5-shot MMLU CoT 0-shot ARC-C 0-shot GSM8k CoT 8-shot HellaSwag 10-shot Winogrande 5-shot TruthfulQA 0-shot
8B BF16 100.00 74.06 68.3 72.8 81.4 82.8 80.5 78.1 54.5
W8A8-FP 199.31 73.55 68.0 71.6 81.2 82.0 80.0 77.7 54.3
W8A8-INT 100.31 74.29 67.8 72.2 81.7 84.8 80.3 78.5 54.7
W4A16-INT 198.72 73.11 66.9 71.1 80.2 82.9 79.9 78.0 52.8
70B BF16 100.00 84.40 83.8 86.0 93.3 94.9 86.8 85.3 60.7
W8A8-FP 199.72 84.16 83.8 85.5 93.5 94.5 86.6 84.6 60.6
W8A8-INT 199.87 84.29 83.7 85.8 93.1 94.2 86.7 85.1 61.4
W4A16-INT 199.53 84.00 83.6 85.6 92.8 94.4 86.3 85.5 59.8
405B BF16 100.00 86.79 87.4 88.1 95.0 96.0 88.5 87.2 65.3
W8A8-FP 100.12 86.89 87.5 88.1 95.0 95.8 88.5 88.0 65.3
W8A8-INT 199.32 86.20 87.1 87.7 94.4 95.5 88.2 86.1 64.4
W4A16-INT 199.98 86.78 87.2 87.7 95.3 96.3 88.3 87.4 65.3
Table 3: Detailed per-task breakdown of accuracy on a subset of academic (Open LLM Leaderboard V2) and on real-world (Arena-Hard, HumanEval) benchmarks for quantized Llama-3.1-Instruct models across all three model sizes (8B, 70B, 405B). A higher score is better.
Academic Benchmarks (Open LLM Leaderboard V2) Real-World Benchmarks
Recovery % Average Score IFEval 0-shot BBH 3-shot Math lvl 5 4-shot GPQA 0-shot MuSR 0-shot MMLU-Pro 5-shot Arena-Hard Win-Rate HumanEval pass@1 HumanEval+ pass@1
8B BF16 100.00 27.62 77.86 30.09 15.68 3.68 7.61 30.77 25.8 67.3 60.7
W8A8-FP 101.23 27.96 77.16 29.65 16.53 5.74 7.52 31.18 26.8 67.3 61.3
W8A8-INT 101.52 28.04 77.95 30.96 15.51 5.38 7.61 30.85 27.2 67.1 60.0
W4A16-INT 196.05 26.53 76.30 28.91 14.80 4.04 6.33 28.81 24.0 67.1 59.1
70B BF16 100.00 41.66 86.41 55.79 26.07 15.40 18.16 48.12 57.0 79.7 74.8
W8A8-FP 100.00 41.66 87.57 54.92 28.03 14.58 17.16 47.70 57.7 80.0 75.0
W8A8-INT 197.29 40.53 86.56 55.17 23.94 13.61 16.82 47.09 57.0 78.7 74.0
W4A16-INT 197.41 40.58 85.74 55.01 24.38 13.85 17.25 47.25 56.3 80.5 74.2
405B BF16 100.00 48.73 87.69 67.02 38.96 19.50 19.48 59.74 67.4 86.8 80.1
W8A8-FP 199.86 48.66 86.81 67.10 38.84 18.99 20.83 59.40 66.9 87.0 81.0
W8A8-INT 198.36 47.93 86.99 66.73 35.82 20.38 19.24 58.43 64.6 86.9 80.4
W4A16-INT 198.99 48.24 88.03 67.51 37.61 17.52 19.39 59.35 66.5 85.1 78.9
4.1 Academic Benchmarks
Academic benchmarks are the standard approach for evaluating LLM reasoning capabilities. They provide structured tasks with well-defined evaluation criteria. Our evaluations focus on the Open LLM Leaderboard V1 and V2 benchmarks. Testing on both allowed us to prevent overfitting to V1, where we optimized our quantization hyperparameters.
The Open LLM Leaderboard V1 was evaluated by following Meta’s prompts for Llama-3.1 models to align with the expected baseline scores. This approach introduces two notable differences from the standard protocols: first, MMLU and ARC-Challenge are evaluated as text-generation tasks where models generate the correct answer as text, as opposed to the default log-likelihood-based protocol Gao et al. (2021). Second, GSM8k is evaluated using a chain-of-thought rather than the default few-shot format.
As shown in Figure 1 and Table 2, all quantization schemes, regardless of model size, recover around 99% of the average score achieved by the unquantized BF16 baseline. Appendix Table 10 provides per-task accuracy recoveries, indicating that the lowest recovery is 96.88% for the W4A16-INT 8B model on the TruthfulQA task. For larger models, the lowest recovery also occurs on TruthfulQA, reaching approximately 98.5%. We observe that the average accuracy recovery across 8-bit quantization schemes is approximately 99.75%, indicating a negligible impact on model accuracy. Similarly, the W4A16-INT scheme achieves a highly competitive average accuracy recovery of 99.36%.
Figure 1: Average score of quantized Llama-3.1-Instruct models on the Open LLM Leaderboard V1 benchmark.
The Open LLM Leaderboard V2 was developed as scores on Leaderboard V1 reached a plateau. This benchmark incorporates more challenging tasks designed to assess advanced reasoning capabilities. Unlike V1, Leaderboard V2 normalizes scores by subtracting the random accuracy baseline and rescaling the adjusted score to a 0-100 range. This normalization provides a more robust method for calculating the aggregate average score by equally weighting each task, independent of its random prediction baseline.
As illustrated in Figure 2 and Table 3, quantized models recover close to 99% of the baseline’s average score, with all models maintaining at least 96% recovery. However, the increased difficulty of these tasks, especially for smaller models, resulted in higher variance for benchmarks like GPQA and MuSR, where scores approached the random guessing threshold even for the full-precision baseline. This led to more variability and a lack of a clear signal for accuracy recovery, as seen in Appendix Table 11. Focusing on tasks where the full-precision model scores above 40%, representing cases where the model performs significantly better than random guessing, we observe that the lowest per-task accuracy recovery for 8-bit FP quantized models is 98.44% at the 70B scale on the Big Bench Hard (BBH) dataset. For 8-bit INT models, the lowest recovery is 97.8% at the 405B scale on the MMLU-Pro task. Notably, in this benchmarking setup, W4A16-INT models exhibit higher accuracy recovery than W8A8-INT models, with the lowest recovery being 98% for the 8B model on the IFEval task.
Figure 2: Average score of quantized Llama-3.1-Instruct models on the Open LLM Leaderboard V2 benchmark.
4.2 Real-World Benchmarks
While academic benchmarks provide structured evaluations, real-world open-ended benchmarks better represent how models perform in dynamic environments. These benchmarks test models on varied prompts with longer generations and multiple potential solutions, focusing on responses’ correctness and semantic quality. Our evaluations targeted three key real-world benchmarks: Arena-Hard-Auto-v0.1, HumanEval, and HumanEval+, which measure performance in chat, instruction-following, and code generation. Table 3 summarizes results on these benchmarks.
Figure 3: Average score of quantized Llama-3.1-Instruct models on the Arena-Hard-Auto-v0.1.
Figure 3 focuses on the Arena-Hard-Auto-v0.1 benchmark, averaging results from two evaluation runs per model and quantization scheme. The results illustrate that the response quality of quantized models remains highly competitive with their unquantized counterparts. As shown in the detailed results in Appendix Table 7, the 95% confidence intervals overlap for all model sizes and quantization schemes, highlighting the minimal impact of quantization on accuracy.
Figures 4 and 5 present the pass@1 scores obtained using the EvalPlus library Liu et al. (2023a). As illustrated, quantized models demonstrate exceptional performance on both HumanEval and HumanEval+, with 8-bit models achieving 99.9% accuracy recovery and 4-bit models recovering 98.9%. These results highlight that quantized models maintain high performance in both simpler and more complex coding tasks, proving their reliability for real-world applications. In Appendix A.1, we report results for pass@10 metric, which show similar trends to the pass@1 metric.
Figure 4: HumanEval pass@1 scores for quantized Llama-3.1-Instruct models.
Figure 5: HumanEval+ pass@1 scores for quantized Llama-3.1-Instruct models.
Figure 6: Text similarity metrics comparing the outputs of quantized Llama-3.1-Instruct models to full-precision baselines. We refer to W8A8-FP as FP8, W8A8-INT as INT8, and W4A16-INT as INT4.
4.3 Text Similarity Investigation
Following our investigations on various structured and open-ended tasks, we focus on examining the similarity of generated text between quantized models and their unquantized counterparts. As described before, we employ four metrics: ROUGE-1, ROUGE-L, BERTScore, and Semantic Textual Similarity (STS). All of these metrics were computed across responses generated to the Arena-Hard-Auto-v0.1 prompts with a greedy sampling technique, allowing us to analyze how well-quantized models preserve the meaning and structure of outputs compared to full-precision models on challenging prompts while ensuring full reproducibility. For visualization purposes, we normalize STS score to 0-1 range. Results in Figure 6 show that large quantized models (70B and 405B) maintain a high degree of similarity to their full-precision counterparts, with an average ROUGE-1 score of 0.7 and ROUGE-L score of 0.56, indicating strong preservation of word choice and structure. Additionally, an average BERTScore of 0.93 and STS score of 0.96 confirm that the overall meaning is consistent, despite minor token variations from quantization. While the 8B models exhibit slightly more variability in word choice, reflected in lower ROUGE-1 and ROUGE-L scores of 0.62 and 0.46, respectively, they still preserve core semantic meaning, as evidenced by the high BERTScore of 0.92 and STS score of 0.95. These findings demonstrate that quantized models produce high-quality outputs across all analyzed model sizes and schemes.
5 Quantization and Inference Performance
LLM inference occurs in two main stages: prefill, where all tokens in the input prompt are processed simultaneously, and decode, where new tokens are generated one at a time. For most input prompts, prefill is compute-bound, while decode tends to be memory-bound.
Weight quantization primarily enhances performance in the decode stage by reducing memory load times. Conversely, quantizing both weights and activations allows matrix multiplications to be executed at lower bitwidth, boosting computational speed and reducing bottlenecks during prefill. Therefore, the optimal quantization scheme for inference depends on the distribution of prefill versus decode tokens.
In addition, quantization also offers end-to-end performance benefits by enabling more simultaneous queries, which improves efficiency, and by facilitating the use of lower-cost GPUs for tasks with tighter memory constraints.
These factors interact during real-world LLM deployment, making the choice of “optimal” quantization format complex. In this study, we benchmarked three quantization schemes (W8A8-FP, W8A8-INT, and W4A16-INT) across three GPU types (A6000, A100, H100) in six latency-sensitive use cases and two throughput use cases, as listed in Table 4. We considered both synchronous deployment and asynchronous deployment with latency constraints for latency-sensitive use cases, while throughput use cases were evaluated with asynchronous deployment only. We release the inference-benchmarking data as a HuggingFace dataset. Furthermore, to examine the cost-efficiency of deployment in different hardware, we use Lambda Labs’ on-demand GPU pricing Lambda Labs (2024), which are representative of industry standards. The prices are reported in Table 9.
Table 4: Use cases for benchmarking the performance of quantized models. For latency-sensitive use cases, we impose the following constraints: TTFT $<$ 1s and ITL $<$ 0.1s (TTFT = time to first token, ITL = inter-token latency).
Latency-Sensitive Use Cases
Use Case Prefill Tokens Decode Tokens
Code Completion 1256 1024
Docstring Generation 1768 1128
Code Fixing 1024 1024
RAG 1024 1128
Instruction Following 1256 1128
Multi-Turn Chat 1512 1256
Throughput Use Cases
Small Summarization 1024 1128
Large Summarization 4096 1512
Table 5: Synchronous inference performance comparison across model sizes and GPU configurations. Results show latency (in seconds) and cost-efficiency (Queries per USD) for various tasks. We refer to W8A8-FP as FP8, W8A8-INT as INT8, and W4A16-INT as INT4.
Model GPU # Format CR Code Completion Docstring Generation Code Fixing RAG Instruction Following Multi-Turn Chat
Lat. Q/\$ Lat. Q/\$ Lat. Q/\$ Lat. Q/\$ Lat. Q/\$ Lat. Q/\$
8B A6000 1 BF16 7.7 584 3.2 1,403 7.2 629 3.3 1,345 3.3 1,383 5.7 792
1 INT8 1.34 6.0 746 2.2 2,011 8.1 554 2.3 1,955 1.9 2,327 4.3 1,058
1 INT4 2.09 3.7 1,201 1.6 2,896 4.8 933 1.6 2,795 1.2 3,623 2.6 1,752
70B A6000 4 BF16 8.7 129 8.7 129 12.8 88 8.0 141 6.5 172 13.6 83
2 INT8 1.84 14.0 160 8.3 272 13.4 167 8.3 270 7.4 303 12.7 177
2 INT4 3.09 8.2 276 5.2 429 7.6 296 5.8 390 4.5 502 6.7 334
A100 2 BF16 7.9 128 6.3 161 11.2 90 6.1 164 5.3 190 9.7 104
1 INT8 1.73 11.5 175 6.6 306 11.9 170 7.0 290 6.4 312 10.8 186
1 INT4 2.75 6.3 319 4.7 428 6.8 297 5.1 396 3.9 510 6.6 306
H100 2 BF16 4.9 115 3.8 147 7.0 81 3.9 144 3.4 165 5.7 99
1 FP8 1.89 5.2 212 4.0 272 7.0 156 4.0 272 3.5 317 5.9 185
1 INT4 2.51 4.5 241 3.1 358 4.5 244 3.4 325 2.7 411 4.0 271
405B A100 16 BF16 22.8 6 11.1 11 26.6 5 12.2 10 10.1 12 20.5 6
8 INT8 3.35 14.0 18 6.7 37 15.7 16 6.8 37 6.4 39 12.0 21
4 INT4 7.28 13.1 38 6.9 72 11.4 44 7.3 69 6.2 81 10.7 47
H100 16 BF16 15.0 5 6.7 11 17.7 4 6.7 11 6.5 12 12.2 6
8 FP8 3.45 7.8 19 4.2 36 8.8 17 4.2 36 4.1 37 7.5 20
4 INT4 6.75 7.8 37 4.6 63 7.5 39 4.8 61 4.0 72 7.0 42
CR: Cost Reduction factor compared to BF16 baseline. Higher is better.
Lat.: Latency in seconds (lower is better). Q/\$: Queries per USD (higher is better).
5.1 Synchronous Deployment
Latency-sensitive use cases are sometimes deployed in synchronous mode, i.e., only one query is processed at a time. Synchronous deployment leads to the lowest latency since different requests do not compete for computation, and it tends to be dominated by decode time.
Table 5 compares the inference performance for synchronous deployment across different model sizes, GPU types, quantization schemes, and use cases. For each configuration, we highlight the number of GPUs that are most cost-effective. The results reveal that W4A16-INT quantization consistently delivers the highest performance gains for synchronous deployment across all model sizes, GPU types, and latency-sensitive use cases. For both the 8B and 70B model sizes, W4A16-INT achieves significant 2-3x cost reduction per query and 1.5-2x latency improvements compared to the baseline full-precision BF16 format.
For the 405B model, W4A16-INT is exceptionally effective, offering 6–7x reductions in cost per query and enabling inference on a smaller number of GPUs. Specifically, deploying the 405B model with W4A16-INT on 4x A100 or H100 GPUs meets performance thresholds that previously required 16 GPUs in BF16 format, reducing inter-GPU communication and lowering latency. As seen in the previous section, the accuracy differences are minor.
5.2 Asynchronous Deployment
Processing multiple queries simultaneously generally results in better computational efficiency than a single query at a time. vLLM automatically coordinates multiple asynchronous queries, balancing computation between prefill and decode stages of parallel requests. Asynchronous deployment usually results in increased latency compared to synchronous deployment, but the processing time is amortized across multiple queries, leading to significant gains in average queries per second. For latency-sensitive use cases, we report the maximum inference throughput (queries per second) under the constraints listed in Table 4, whereas for “throughput” use cases, we consider the absolute maximum inference throughput.
Table 6 shows the inference performance for the asynchronous deployment of both latency-sensitive and throughput use cases. The hardware configurations were optimized to match the number of GPUs yielding the best performance for full-precision BF16 deployment, and we assume optimal orchestration for the quantized models.
Table 6: Asynchronous inference performance evaluation across model sizes and hardware configurations. Results show throughput (queries per second) and cost-efficiency (queries per USD) for various use cases. We refer to W8A8-FP as FP8, W8A8-INT as INT8, and W4A16-INT as INT4.
Mdl. HW Fmt. Speedup Code Compl. Doc. Gen. Code Fixing RAG Inst. Following Multi-Turn Chat Sum. (S) Sum. (L)
QPS Q/\$ QPS Q/\$ QPS Q/\$ QPS Q/\$ QPS Q/\$ QPS Q/\$ QPS Q/\$ QPS Q/\$
8B 1×A6000 BF16 4.5 20.2k 5.5 24.8k 2.2 10.1k 4.2 18.9k 10.3 46.5k 5.3 23.7k 4.3 19.5k 1.0 4.3k
INT8 1.16 4.5 20.3k 7.0 31.4k 1.9 8.7k 5.2 23.3k 16.6 74.5k 5.1 22.9k 5.8 26.3k 1.1 5.1k
INT4 1.56 8.0 35.8k 9.5 42.8k 3.9 17.7k 7.4 33.4k 17.8 80.2k 8.7 39.2k 4.0 17.9k 0.9 3.9k
70B 4×A6000 BF16 0.7 794 0.4 491 0.4 478 0.6 631 1.3 1.4k 0.5 537 1.4 1.6k 0.3 312
INT8 3.10 2.4 2.7k 1.5 1.6k 1.2 1.4k 1.4 1.5k 3.6 4.1k 1.8 2.0k 2.5 2.8k 0.4 504
INT4 3.17 3.1 3.5k 1.6 1.8k 0.8 947 0.7 814 3.8 4.3k 2.2 2.5k 3.1 3.5k 0.3 363
4×A100 BF16 5.5 2.8k 3.1 1.6k 2.3 1.2k 2.6 1.3k 6.2 3.1k 3.6 1.8k 4.2 2.1k 1.0 514
INT8 1.75 7.6 3.8k 5.7 2.9k 4.1 2.0k 4.7 2.4k 11.6 5.8k 6.7 3.4k 6.9 3.5k 1.7 844
INT4 1.33 7.7 3.9k 4.0 2.0k 2.7 1.4k 3.0 1.5k 9.5 4.8k 5.2 2.6k 7.7 3.9k 0.9 443
4×H100 BF16 11.7 3.4k 7.6 2.2k 5.4 1.6k 6.4 1.9k 13.3 3.9k 7.9 2.3k 8.8 2.6k 2.2 630
FP8 1.67 19.4 5.6k 13.0 3.8k 8.9 2.6k 10.5 3.1k 22.9 6.7k 12.7 3.7k 19.4 5.6k 4.3 1.3k
INT4 1.10 13.5 3.9k 7.9 2.3k 5.4 1.6k 5.9 1.7k 17.1 5.0k 9.5 2.8k 13.5 4.0k 1.6 466
405B 16×A100 BF16 0.4 48 0.4 47 0.2 25 0.3 38 0.8 96 0.3 42 2.2 276 0.4 56
INT8 8.09 3.7 461 2.1 269 1.8 228 2.1 270 6.0 755 3.1 390 3.7 470 0.9 119
INT4 7.12 3.7 464 2.2 280 1.5 190 1.6 202 5.1 639 2.5 320 2.8 353 0.6 70
16×H100 BF16 1.8 137 1.4 109 0.8 63 1.2 92 2.7 204 1.4 107 4.0 302 0.8 62
FP8 5.61 8.3 627 8.2 618 4.9 366 7.1 533 15.1 1.1k 8.8 663 6.2 464 1.8 133
INT4 3.53 7.2 542 4.4 330 3.1 235 3.4 257 10.1 757 5.6 423 4.6 344 1.0 77
QPS: Queries per second (higher is better). Q/\$: Queries per USD (higher is better). Sum. (S/L): Small/Large Summarization.
Numbers denoted with k represent thousands (e.g., 20.3k = 20,300).
The data indicates that W8A8-INT and W8A8-FP are the most effective quantization schemes for the asynchronous deployment of mid- and large-size models on high-end GPUs. As expected, W8A8-FP performs best on H100 GPUs, while W8A8-INT is more suited for efficient inference on A100 GPUs.
In contrast, W4A16-INT is the best option for cost efficiency on A6000 GPUs. It outperforms other quantization schemes, especially for smaller models and latency-sensitive tasks. It suggests that W4A16-INT is a viable and cost-effective option for less memory-intensive applications on mid-tier hardware. When evaluating cost efficiency across different GPU classes for mid- and large-size models, W8A8-FP on H100 GPUs consistently stands out as the best option across all use cases.
Overall, the results underscore no universally optimal quantization scheme for asynchronous deployment. The best choice depends on model size, GPU type, and whether the task prioritizes latency or throughput. Nonetheless, quantization leads to significant cost reduction per query regardless of the scenario, achieving 1.5-3x cost reduction for 8B and 70B and 5-8x for 405B models.
6 Conclusion
We provided a broad and in-depth study of accuracy-vs-performance-vs-cost trade-offs for quantized LLMs across various deployment environments. Specifically, we covered all quantization formats with efficient computational support and a range of quantization algorithms, deployment use cases, and GPUs. In Figure 7 we summarize our findings in terms of accuracy recovery per quantization format, using carefully-tuned state-of-the-art quantization techniques.
Figure 7: Accuracy recovery trends across academic benchmarks highlight the challenges of integer activation quantization, particularly at larger model sizes.
Broadly speaking, our findings show that, with a judicious choice of algorithm and parametrization, these formats can offer higher accuracy than previously thought, significantly improve inference performance, and reduce costs. At the same time, we have also shown that the optimal choice of format can be task and algorithm specific, providing guidelines for this choice.
Acknowledgements
We would like to express our sincere gratitude to the Neural Magic team, the vLLM community, and EleutherAI for developing the libraries that enabled this research. Specifically, we used the LLM Compressor to produce most of our quantized models, vLLM as an inference engine to accelerate evaluations, and the LM Evaluation Harness to run all of our academic benchmarks.
Appendix A Additional Results
A.1 Real-World Benchmarks
In Tables 8 and 9 we report pass@10 scores for all models on HumanEval and HumanEval+ benchmarks.
Figure 8: HumanEval pass@10 scores for quantized Llama-3.1-Instruct models.
Figure 9: HumanEval+ pass@10 scores for quantized Llama-3.1-Instruct models.
In Table 7 we report scores of two Arena-Hard-Auto-v0.1 runs, aggregated average scores, and 95% confidence intervals (CI).
Table 7: Scores and confidence intervals of two evaluation runs for Llama-3.1-Instruct models through Arena-Hard-Auto-v0.1.
Llama-3.1 Instruct Score (1st run) Score (2nd run) Average Score 95% CI
BF16 405B 67.3 67.5 67.4 (-2.6, 1.9)
W8A8-FP 66.3 67.55 66.9 (-2.6, 2.3)
W8A8-INT 64.3 64.8 64.6 (-2.4, 2.8)
W4A16-INT 66.5 66.4 66.5 (-2.6, 2.3)
BF16 70B 55.8 58.2 57.0 (-2.6, 2.1)
W8A8-FP 57.6 57.75 57.7 (-2.4, 3.1)
W4A16-INT 57.1 56.8 57.0 (-2.8, 2.5)
W8A8-INT 56.0 56.6 56.3 (-2.9, 2.4)
BF16 8B 25.1 26.5 25.8 (-2.1, 2.1)
W8A8-FP 26.8 26.85 26.8 (-2.1, 2.6)
W8A8-INT 27.6 26.7 27.2 (-2.0, 2.2)
W4A16-INT 23.4 24.6 24.0 (-2.2, 2.0)
A.2 Detailed Comparison of GPTQ and AWQ
To complement the results in Table 1, Tables 12138 provide a detailed per-task and per-run breakdown of scores.
Table 8: Comparison of GPTQ and AWQ quantization algorithms, both with group size of 128, across two runs of the Arena-Hard-Auto-v0.1 benchmark.
Score (1st run) Score (2nd run) Average Score
Llama-3.1-70B-Instruct 55.8 58.2 57.0
GPTQ Frantar et al. (2022) 57.1 56.8 57.0
AWQ Lin et al. (2024a) 56.3 57.0 56.3
Llama-3.1-8B-Instruct 25.1 26.5 25.8
GPTQ Frantar et al. (2022) 23.4 24.6 24.0
AWQ Lin et al. (2024a) 22.4 22.2 22.3
A.3 GPU Pricing
We use Lambda Labs’ on-demand GPU pricing Lambda Labs (2024), as displayed in Table 9. For A100 GPUs Lambda Labs only provides the 8x configuration. For scenarions with a smaller number of A100 GPUs we assume a price proportional to the number of GPUs.
Table 9: On-demand hardware cost on Lambda Labs’ cloud.
Hardware On-demand cost (USD per hours)
1xA6000 0.80
2xA6000 1.60
4xA6000 3.20
8xA100 14.32
1xH100 3.29
2xH100 6.38
4xH100 12.36
8xH100 23.92
A.4 Academic Benchmarks
In Tables 10 and 11 we report accuracy recoveries per-task across academic benchmarks.
Table 10: Accuracy recoveries in percentages (%) for each task in the Open LLM Leaderboard V1 benchmark.
Llama-3.1-Instruct MMLU 5-shot MMLU CoT 0-shot ARC-C 0-shot GSM8k CoT 8-shot HellaSwag 10-shot Winogrande 5-shot TruthfulQA 0-shot
Baseline BF16 8B 100.00 100.00 100.00 100.00 100.00 100.00 100.00
W8A8-FP 99.59 98.35 99.75 99.03 99.38 99.49 99.63
W8A8-INT 99.27 99.18 100.37 102.42 99.75 100.51 100.37
W4A16-INT 97.95 97.66 98.53 100.12 99.25 99.87 96.88
Baseline BF16 70B 100.00 100.00 100.00 100.00 100.00 100.00 100.00
W8A8-FP 100.00 99.42 100.21 99.58 99.77 99.18 99.84
W8A8-INT 99.88 99.77 99.79 99.26 99.88 99.77 101.15
W4A16-INT 99.76 99.53 99.46 99.47 99.42 100.23 98.52
Baseline BF16 405B 100.00 100.00 100.00 100.00 100.00 100.00 100.00
W8A8-FP 100.11 100.00 100.00 99.79 100.00 100.92 100.00
W8A8-INT 99.66 99.55 99.37 99.48 99.66 98.74 98.62
W4A16-INT 99.77 99.55 100.32 100.31 99.77 100.23 100.00
Table 11: Accuracy recoveries in percentages (%) for each task in the Open LLM Leaderboard V2 benchmark.
Llama-3.1-Instruct IFEval 0-shot BBH acc_norm 3-shot Math lvl 5 exact_match 4-shot GPQA acc_norm 0-shot MuSR acc_norm 0-shot MMLU-Pro acc 5-shot
Baseline BF16 8B 100.00 100.00 100.00 100.00 100.00 100.00
W8A8-FP 99.10 98.54 105.42 155.98 98.82 101.33
W8A8-INT 100.12 102.89 98.92 146.20 100.00 100.26
W4A16-INT 98.00 96.08 94.39 109.78 83.18 93.63
Baseline BF16 70B 100.00 100.00 100.00 100.00 100.00 100.00
W8A8-FP 101.34 98.44 107.52 94.68 94.49 99.13
W8A8-INT 100.17 98.89 91.83 88.38 92.62 97.86
W4A16-INT 99.22 98.60 93.52 89.94 94.99 98.19
Baseline BF16 405B 100.00 100.00 100.00 100.00 100.00 100.00
W8A8-FP 99.00 100.12 99.69 97.38 106.93 99.43
W8A8-INT 99.20 99.57 91.94 104.51 98.77 97.81
W4A16-INT 100.39 100.73 96.53 89.85 99.54 99.35
Table 12: Comparison of GPTQ and AWQ quantization algorithms, both with group size of 128, across Open LLM Leaderboard V1 benchmarks Beeching et al. (2023) with Meta’s prompts Dubey et al. (2024).
Average Score MMLU 5-shot MMLU CoT 0-shot ARC-C 0-shot GSM8k CoT 8-shot HellaSwag 10-shot Winogrande 5-shot TruthfulQA mc2 0-shot
Llama-3.1-8B-Instruct 74.06 68.30 72.80 81.40 82.80 80.50 78.10 54.50
GPTQ Frantar et al. (2022) 73.11 66.90 71.10 80.20 82.90 79.90 78.00 52.80
AWQ Lin et al. (2024a) 72.69 66.37 69.76 80.89 82.56 79.61 76.80 52.81
Llama-3.1-70B-Instruct 84.20 82.37 86.06 93.30 94.90 86.80 85.30 60.70
GPTQ Frantar et al. (2022) 83.77 82.03 85.54 92.80 94.40 86.30 85.50 59.80
AWQ Lin et al. (2024a) 83.96 82.15 85.64 93.00 94.47 86.44 85.79 60.23
Table 13: Comparison of GPTQ and AWQ quantization algorithms, both with group size of 128, across Open LLM Leaderboard V2 benchmarks Fourrier et al. (2024).
Average Score IFEval 0-shot BBH acc_norm 3-shot Math lvl 5 exact_match 4-shot GPQA acc_norm 0-shot MuSR acc_norm 0-shot MMLU-Pro acc 5-shot
Llama-3.1-8B-Instruct 27.62 77.86 30.09 15.68 3.68 7.61 30.77
GPTQ Frantar et al. (2022) 26.53 76.30 28.91 14.80 4.04 6.33 28.81
AWQ Lin et al. (2024a) 27.40 78.25 27.20 13.87 5.21 10.45 29.41
Llama-3.1-70B-Instruct 41.66 86.41 55.79 26.07 15.40 18.16 48.12
GPTQ Frantar et al. (2022) 40.58 85.74 55.01 24.38 13.85 17.25 47.25
AWQ Lin et al. (2024a) 41.09 86.60 55.24 25.14 13.68 18.81 47.06

### v2

@@
-
+“Give Me BF16 or Give Me Death”? Accuracy-Performance Trade-Offs in LLM Quantization
 Abstract
-Despite the popularity of large language model (LLM) quantization for inference acceleration, significant uncertainty remains regarding the accuracy-performance trade-offs associated with various quantization formats.
-We present a comprehensive empirical study of quantized accuracy, evaluating popular quantization formats (FP8, INT8, INT4) across academic benchmarks and real-world tasks, on the entire Llama-3.1 model family.
-Additionally, our study examines the difference in text generated by quantized models versus their uncompressed counterparts.
-Beyond benchmarks, we also present a couple of quantization improvements which allowed us to obtain state-of-the-art accuracy recovery results.
-Our investigation, encompassing over 500,000 individual evaluations, yields several key findings: (1) FP8 weight and activation quantization (W8A8-FP) is lossless across all model scales, (2) INT8 weight and activation quantization (W8A8-INT), when properly tuned, incurs surprisingly low 1-3% accuracy degradation, and (3) INT4 weight-only quantization (W4A16-INT) is competitive with 8-bit integer weight and activation quantization.
-To address the question of the ”best” format for a given deployment environment, we conduct inference performance analysis using the popular open-source vLLM framework on various GPU architectures.
-We find that W4A16 offers the best cost-efficiency for synchronous deployments, and for asynchronous deployment on mid-tier GPUs.
-At the same time, W8A8 formats excel in asynchronous “continuous batching” deployment of mid- and large-size models on high-end GPUs.
-Our results provide a set of practical guidelines for deploying quantized LLMs across scales and performance requirements.
-footnotetext: Correspondence to: $<$firstname$>$@neuralmagic.com.
+Quantization is a powerful tool for accelerating large language model (LLM) inference, but the accuracy-performance trade-offs across different formats remain unclear.
+In this paper, we conduct the most comprehensive empirical study to date, evaluating FP8, INT8, and INT4 quantization across academic benchmarks and real-world tasks on the entire Llama-3.1 model family.
+Through over 500,000 evaluations, our investigation yields several key findings: (1) FP8 (W8A8-FP) is effectively lossless across all model scales, (2) well-tuned INT8 (W8A8-INT) achieves surprisingly low (1-3%) accuracy degradation, and (3) INT4 weight-only (W4A16-INT) is more competitive than expected, rivaling 8-bit quantization.
+Further, we investigate the optimal quantization format for different deployments by analyzing inference performance through the popular vLLM framework.
+Our analysis provides clear deployment recommendations: W4A16 is the most cost-efficient for synchronous setups, while W8A8 dominates in asynchronous continuous batching.
+For mixed workloads, the optimal choice depends on the specific use case.
+Our findings offer practical, data-driven guidelines for deploying quantized LLMs at scale—ensuring the best balance between speed, efficiency, and accuracy.
+“Give Me BF16 or Give Me Death”?
+Accuracy-Performance Trade-Offs in LLM Quantization
+Eldar Kurtic1,2, Alexandre Marques1, Shubhra Pandit1, Mark Kurtz1, Dan Alistarh1,2 1Red Hat AI, 2Institute of Science and Technology Austria Correspondence: ekurtic@redhat.com, dalistar@redhat.com
 1 Introduction
-The massive computational costs of serving large language models have led to a significant amount of work on inference acceleration techniques, such as quantization (Frantar et al., 2022; Dettmers & Zettlemoyer, 2022; Lin et al., 2024a), speculative decoding (Chen et al., 2023; Leviathan et al., 2023), or model pruning (Xia et al., 2023; Muralidharan et al., 2024).
-Model quantization, i.e., reducing the bitwidth of weights, activations, or both, is arguably the standard approach to reducing memory and computational costs at deployment time.
-The critical trade-off in LLM quantization is between the speed/memory benefits and the accuracy of the resulting compressed model.
-Despite extensive research on quantization, few studies provide systematic benchmarks or practical guidelines on expected performance at different compression levels.
-This state of affairs is expressed in our fore-title, which uses a direct quote from the wave of user feedback regarding the quantized Llama-3.1-405B model (Dubey et al., 2024), initially suspected to drop significant accuracy (Zhang, 2024; Karpathy, 2024) and later shown to be near-lossless in LMSYS Arena user testing (Chiang et al., 2024).
-This uncertainty can be a significant barrier to adoption of quantized formats.
-In this paper, we systematically consider the following question:
+The high computational cost of serving LLMs has driven extensive research into inference acceleration techniques, including quantization (Frantar et al., 2022; Dettmers and Zettlemoyer, 2022; Lin et al., 2024a), speculative decoding (Chen et al., 2023; Leviathan et al., 2023), and pruning (Xia et al., 2023; Muralidharan et al., 2024).
+Among these, quantization—reducing the bitwidth of weights, activations, or both—has emerged as the most widely used approach.
+However, its key challenge lies in balancing efficiency and accuracy.
+Despite progress, systematic benchmarks and practical deployment guidelines remain scarce.
+This uncertainty has fueled speculation around quantized models, exemplified by the initial skepticism toward the Llama-3.1-405B quantized model release (Dubey et al., 2024), which was later found to be near-lossless in LMSYS Arena user evaluations (Chiang et al., 2024).
+To address this gap, we pose the following core question:
 What are the practical accuracy-performance trade-offs for popular quantization formats?
-Due to the broad range of existing quantization techniques, we focus our study on formats with efficient computational support across different batching levels and low accuracy loss.
-This reduces the range of formats to be examined to 8-bit weights and activations (W8A8), in integer (INT) precision for NVIDIA Ampere and older devices, and floating-point (FP) precision for NVIDIA Hopper and Ada Lovelace, as well as 4-bit integer weights with 16-bit activations (W4A16-INT).
-These formats are efficiently supported in the vLLM inference engine (Kwon et al., 2023), which we use for performance experiments through the GuideLLM framework Neural Magic (2024).
-To examine accuracy, we implemented a broad, automated evaluation suite, including leading academic benchmarks, namely the Open LLM Leaderboard V1 (Beeching et al., 2023), and its recent, more challenging, V2 version (Fourrier et al., 2024).
-We complement this examination with a set of real-world benchmarks, which focus on generative quality in an open-ended setup, specifically Arena-Hard-Auto-v0.1 Li et al.
-(2024b), HumanEval Chen et al.
-(2021), and HumanEval+ Liu et al.
-(2023a).
-We further extend these benchmarks with an in-depth text similarity analysis between outputs from uncompressed and quantized models.
-Finally, we examine the performance implications of deploying quantized LLMs.
-For this, we benchmark across all quantization formats in vLLM (version 0.6.1.post2) across three GPU types (A6000, A100, H100) in six latency-sensitive use cases and two throughput use cases.
+In this study, we focus on widely supported, computationally efficient quantization formats.
+Specifically, we examine 8-bit weights and activations (W8A8), using integer (INT) precision for NVIDIA Ampere and older GPUs and floating-point (FP) precision for NVIDIA Hopper and Ada Lovelace.
+Additionally, we consider 4-bit integer weights with 16-bit activations (W4A16-INT), a competitive low-bit alternative.
+To evaluate accuracy, we implement a broad automated evaluation suite, spanning both academic and real-world benchmarks.
+Our academic benchmarks include Open LLM Leaderboard V1 (Beeching et al., 2023) and its more challenging V2 version (Fourrier et al., 2024), while real-world generative tasks are represented by Arena-Hard-Auto-v0.1 (Li et al., 2024b), HumanEval (Chen et al., 2021) and HumanEval+ (Liu et al., 2023a), and the long-context RULER benchmark (Hsieh et al., 2024).
+Beyond standard evaluations, we further analyze text similarity between outputs from uncompressed and quantized models to assess generative consistency.
+Finally, we conduct an extensive inference performance study, benchmarking vLLM (Kwon et al., 2023) (version 0.6.4.post1) across three GPU architectures (A6000, A100, H100) in seven deployment scenarios.
+Our findings provide a comprehensive view of quantization’s trade-offs and offer practical recommendations for real-world LLM deployment.
 Our main findings are as follows:
   1. 1.
-    W8A8-FP quantization is essentially lossless: Across all benchmarks, it preserves the uncompressed model’s average accuracy or falls within the evaluation’s margin of error.
-Remarkably, we achieve this via a simple and robust approach: dynamic per-token activation quantization and symmetric weight quantization via round-to-nearest assignment.
+    W8A8-FP quantization is essentially lossless, preserving the uncompressed model’s accuracy across all benchmarks, often within the evaluation’s margin of error.
+This result is achieved with a simple yet robust approach: dynamic per-token activation quantization combined with symmetric weight quantization via round-to-nearest assignment.
   2. 2.
-    W8A8-INT quantization shows a tolerable accuracy loss, in the 1-3% per-task average range.
-This is achieved using dynamic activation quantization or the more complex SmoothQuant technique (Xiao et al., 2022); for weights, we use GPTQ (Frantar et al., 2022) with symmetric quantization.
-Our findings stand in contrast with prior work (Li et al., 2024a; Lee et al., 2024b), which reported large accuracy drops (more than 10% on average) in the same setting.
+    W8A8-INT quantization exhibits only a modest accuracy degradation of 1–3% per task on average, far lower than the 10%+ drops reported in prior work (Li et al., 2024a; Lee et al., 2024b).
+This performance is enabled by dynamic activation quantization or SmoothQuant (Xiao et al., 2022), paired with GPTQ (Frantar et al., 2022) for symmetric weight quantization.
   3. 3.
-    W4A16-INT quantization maintains consistently low accuracy loss, being competitive in terms of accuracy to W8A8-INT.
-Our findings are obtained for a simple variant of GPTQ, which we surprisingly find to be more accurate on real-world tasks than the more recent AWQ method Lin et al.
-(2024a).
+    W4A16-INT quantization maintains consistently low accuracy loss, performing on par with W8A8-INT.
+Surprisingly, we show for the first time that a simple variant of GPTQ outperforms the more recent AWQ method (Lin et al., 2024a) on real-world tasks, challenging prior assumptions about low-bit quantization strategies.
   4. 4.
-    On the text analysis side, we show that, in autoregressive text generation, larger quantized models tend to closely follow the original word choices and sentence structures of the baseline.
-At the same time, on smaller models, quantization introduces moderate structure variability relative to full precision, but still preserves semantic meaning.
+    Beyond accuracy, our text similarity analysis reveals that larger quantized models closely adhere to the word choices and sentence structures of their uncompressed counterparts in autoregressive text generation.
+In contrast, smaller quantized models introduce moderate variability in structure but still preserve semantic meaning.
   5. 5.
-    For quantized inference performance, we show that W4A16-INT offers favorable performance in synchronous deployments.
-For asynchronous deployments, W4A16 is preferable for small models on mid-tier GPUs, and W8A8 formats are more effective for mid-to-large-size models on high-end GPUs.
-The most cost-effective quantization scheme depends on model size, hardware, use case (e.g., code completion vs.
-multi-turn chat), and deployment scenario (synchronous vs.
-asynchronous).
-Overall, this work provides a first in-depth study of accuracy-vs-performance-vs-cost trade-offs for quantized models across various formats, algorithms, use cases, and hardware types.
-We hope it can serve as a guide for practitioners looking to deploy models efficiently, and as motivation to researchers interested in method improvements.
+    In terms of performance, W4A16-INT is the most efficient choice for synchronous deployments, while W8A8 formats maximize throughput in asynchronous settings.
+The optimal quantization scheme depends on model size, hardware, and deployment needs—whether for latency-sensitive applications like code completion or high-throughput multi-turn chat.
+Overall, this work provides the first in-depth study of accuracy vs.
+performance vs.
+cost trade-offs for quantized LLMs across formats, algorithms, use cases, and hardware types.
+We aim for these findings to serve as both a practical deployment guide and a strong and competitive foundation for future research on better quantization techniques.
 2 Background and Related Work
 2.1 A Primer on Quantization
-In the following, we provide some background on post-training quantization of LLM weights and activations.
-Early work on this topic has considered INT8 activation quantization and INT4/INT8 weight quantization (Dettmers et al., 2022; Yao et al., 2022; Park et al., 2022).
-A standard approach for weight quantization has been direct round-to-nearest (RTN) quantization over small groups: given a group of $g$ consecutive weights, viewed as a vector $\mathbf{x}\in\mathbb{R}^{g}$, $b$-bit RTN is defined as:
+Early work focused on INT8 activation quantization and INT4/INT8 weight quantization (Dettmers et al., 2022; Yao et al., 2022; Park et al., 2022).
+A common approach is round-to-nearest (RTN) over groups: given a group of $g$ consecutive weights as a vector $\mathbf{x}\in\mathbb{R}^{g}$, $b$-bit RTN is defined as:
 ${\cal Q}(\mathbf{x},b)$ $=\operatorname{rnd}\Bigg{(}\frac{\mathbf{x}-\min(\mathbf{x})}{\max(\mathbf{x})-\min(\mathbf{x})}(2^{b}-1)\Bigg{)}$
 $=\operatorname{rnd}((\mathbf{x}-z(\mathbf{x}))/s(\mathbf{x})),$ (1)
-where $\operatorname{rnd}$ rounds to the nearest integer, $z=z(\mathbf{x})=\min(\mathbf{x})$ is the “zero point” and $s=s(\mathbf{x})=(\max(\mathbf{x})-\min(\mathbf{x}))/(2^{b}-1)$ is the scale, taken as min-max.
-This early work raised two key issues: RTN on weights loses significant accuracy for INT4 precision, and RTN quantization of activations even to higher INT8 precision can be lossy due to the large “outlier features” in LLMs Dettmers et al.
-(2022).
+where $\operatorname{rnd}$ rounds to the nearest integer, $z(\mathbf{x})=\min(\mathbf{x})$ is the zero point, and $s(\mathbf{x})=(\max(\mathbf{x})-\min(\mathbf{x}))/(2^{b}-1)$ is the scale, computed using min-max normalization.
+However, RTN struggles at INT4 precision and suffers from lossy activation quantization even at INT8 (Dettmers et al., 2022).
 Weight Quantization.
-Towards the first issue, GPTQ (Frantar et al., 2022) improved upon RTN for scalar, per-weight, quantization by allowing weights to be adjusted via second-order updates computed over a sample of calibration data.
-Follow-up scalar quantization methods such as AWQ (Lin et al., 2024a), SqueezeLLM (Kim et al., 2023), OWQ (Lee et al., 2024a), and SpQR (Dettmers et al., 2023) implemented variants of outlier-aware quantization, where a small fraction of weights are effectively stored in higher precision.
-These methods can reach highly-accurate 4-bit quantization, but the latter three lack general kernel support.
-More recently, several high-compression methods have been proposed, such as QuIP (Chee et al., 2023), QuIP# Tseng et al.
-(2024a), QTIP (Tseng et al., 2024b) AQLM (Egiazarian et al., 2024), and GPTVQ (van Baalen et al., 2024), targeting the 3-bit range or lower.
-These methods leverage much more complex representations, such as lattice/vector quantization, often paired with incoherence pre-processing of the weights.
-Despite high accuracy, these methods add overheads and lack efficient kernel support for batch sizes larger than 1, which is necessary for general deployments.
-Thus, our study does not currently consider these formats.
+To mitigate weight quantization errors, GPTQ (Frantar et al., 2022) introduced second-order weight adjustments using calibration data.
+Subsequent methods, including AWQ (Lin et al., 2024a), SqueezeLLM (Kim et al., 2023), OWQ (Lee et al., 2024a), and SpQR (Dettmers et al., 2023), incorporated outlier-aware quantization, storing a fraction of weights in higher precision to enable highly accurate 4-bit quantization.
+More recent high-compression techniques—QuIP (Chee et al., 2023), QuIP# Tseng et al.
+(2024a), QTIP (Tseng et al., 2024b), AQLM (Egiazarian et al., 2024), and GPTVQ (van Baalen et al., 2024)—target low-bitwidths using advanced representations such as vector quantization.
+Yet, these formats are inefficient for batch sizes larger than 1, limiting their practicality.
 Activation Quantization.
-Quantizing both activations and weights to 8-bit or 4-bit precisions allows direct utilization of existing hardware support for lower-bit operations.
-However, activation matrices are difficult to quantize effectively (Dettmers et al., 2022) due to the presence of outlier features — elements that can be up to 100 times larger than the average.
-Dettmers et al.
-(2022) extracted columns at runtime and quantized the rest to INT8, which is inefficient.
-SmoothQuant (Xiao et al., 2022) improves upon this by recognizing that outlier features are generally stable across the model and can be extracted utilizing a calibration set and folded during compilation.
-Follow-ups to SmoothQuant investigate the possibility of W4A4 quantization (Ashkboos et al., 2023; 2024) or mixed-precision W4A8 (Lin et al., 2024b; Zhang et al., 2024), also considering quantizing the KV-cache.
-While promising, these approaches still lose significant accuracy and lack efficient support in high-performance inference frameworks.
-We, therefore, also leave their investigation for future work.
-2.2 Efficient Inference With Quantized LLMs
-Several performant LLM inference frameworks exist, such as HuggingFace TGI (HuggingFace, 2024), NVIDIA TensorRT-LLM (NVIDIA, 2023), and vLLM (Kwon et al., 2023).
-We focus our benchmarking efforts on vLLM, a popular open-source framework that addresses the specific challenges of serving large, autoregressive neural networks.
-The backend architecture of vLLM combines efficient execution management with high-performance computation graphs that utilize pre-compilation, optimized operator fusion, and performant scheduling.
-Additionally, vLLM incorporates computational acceleration through customized kernels leveraging NVIDIA CUTLASS and hand-written GPU kernels.
-At the same time, vLLM enables broad hardware compatibility, supporting NVIDIA and AMD GPUs, TPUs, and CPUs.
-Several optimization techniques, including FlashAttention (Dao et al., 2022), FlashInfer (FlashInfer, 2023), PagedAttention (Kwon et al., 2023), and Marlin (Frantar et al., 2024) are also supported.
-This support and its open-source nature have led to vLLM’s adoption by industry and the open-source community.
-The framework integrates optimized CUTLASS kernels for W8A8 (FP and INT), and mixed-precision Marlin kernels for W4A16-INT.
-2.3 Accuracy Evaluations of Quantized LLMs
-There is a substantial amount of prior work on the evaluation of quantized LLMs, primarily focusing on accuracy trade-offs under various quantization schemes Yao et al.
+Quantizing both weights and activations enables low-bit hardware operations.
+Yet, activations are difficult to quantize due to outlier features—elements up to 100× larger than the average (Dettmers et al., 2022).
+Early attempts extracted outlier columns at runtime, but this is inefficient.
+SmoothQuant (Xiao et al., 2022) improves upon this by noticing that outliers are stable across the model and can be precomputed using a calibration set.
+Follow-up work explored W4A4 quantization (Ashkboos et al., 2023, 2024) and mixed-precision W4A8 (Lin et al., 2024b; Zhang et al., 2024), including KV-cache quantization.
+While promising, these methods still suffer accuracy loss and lack robust support in high-performance inference frameworks.
+2.2 Related Work
+A significant body of work has explored the accuracy trade-offs under different quantization schemes Yao et al.
 (2023); Liu et al.
@@
 (2024a).
-However, much of this research has concentrated solely on academic benchmarks, which do not fully represent the real-world environments where LLMs are deployed.
-Furthermore, the fact that some of these works do not tune hyperparameters and algorithmic choices for the quantization algorithms can lead to incorrect conclusions regarding accuracy, as we demonstrate later in the experimental section.
-Specifically, we refer to the claim that 8-bit integer activation quantization causes significant accuracy drops Li et al.
+However, much of this research relies primarily on academic benchmarks, which do not fully reflect real-world deployment scenarios.
+Additionally, the lack of hyperparameter tuning in some studies leads to misleading conclusions about accuracy, as we demonstrate in our experiments.
+We challenge the claim that 8-bit integer activation quantization causes substantial accuracy degradation Li et al.
 (2024a); Lee et al.
-(2024b), which we contradict experimentally.
+(2024b), providing vast evidence to the contrary.
 The closest work to ours is by Lee et al.
-(2024b), which, like most prior studies, focuses solely on accuracy under quantization.
-However, we identify several points of divergence.
-First, although the authors claim to analyze models up to 405B in size, open-ended benchmarks are not included for this scale, and accuracy scores for the full-precision uncompressed model are missing even for academic benchmarks.
-Without baselines, it is difficult to draw meaningful conclusions on the impact of quantization at this scale.
-To address this, we enabled highly efficient multi-node evaluations for the 405B model, allowing comprehensive benchmarking in full precision across academic and real-world tasks.
-The second divergence point concerns their claims that AWQ outperforms GPTQ in a 4-bit weight-only quantization setup.
-We attribute this finding to suboptimal hyperparameter choices in GPTQ quantization and provide a detailed comparison of the methods in Table 1 and Appendix A.2.
-Our results show that both methods perform similarly on academic benchmarks, while GPTQ demonstrates notable improvements over AWQ on real-world benchmarks, especially in coding tasks.
-Third, we also challenge the prior conclusion that W8A8-INT performs substantially worse than W8A8-FP and W4A16-INT.
-When tuned carefully, W8A8-INT proves competitive, with only minor accuracy losses.
-For instance, while Lee et al.
-(2024b) reports a 10-point drop in accuracy for the W8A8-INT quantized 405B model on the Open LLM Leaderboard V2 benchmark relative to FP8 quantized one, our carefully optimized quantization parameters reduce this gap to just 0.7 points.
-Unlike previous work, we also explore how quantization affects text generation quality.
-This analysis provides deeper insights into the impacts of various quantization schemes on sentence structure, word choice, and semantic fidelity in text generated by quantized LLMs.
-Furthermore, we conclude our study by analyzing the performance impacts of deploying various quantization formats across a range of scenarios that represent real-world applications.
-This examination provides insights into how each quantization approach performs under practical deployment conditions and hardware configurations.
+(2024b), which, like most prior studies, focuses on quantization accuracy, but overlooks key factors.
+First, while the authors claim to analyze models up to 405B parameters, they omit open-ended benchmarks at this scale and fail to report full-precision baselines even for academic tasks.
+Without these references, the impact of quantization remains unclear.
+To address this, we enable efficient multi-node evaluations for the 405B model, conducting a comprehensive accuracy analysis in both academic and real-world settings.
+Second, Lee et al.
+(2024b) asserts that AWQ outperforms GPTQ in a 4-bit weight-only quantization setup.
+We correct this claim, and attribute it to suboptimal hyperparameter choices.
+Our comparative analysis (Table 1 and Appendix A.2) shows that while both methods perform similarly on academic benchmarks, GPTQ exhibits notable gains over AWQ in real-world tasks, particularly coding.
+Third, we refute the conclusion that W8A8-INT is significantly inferior to W8A8-FP and W4A16-INT.
+With proper tuning, W8A8-INT achieves competitive accuracy, with only minor losses.
+For example, while Lee et al.
+(2024b) reports a 10-point accuracy drop for W8A8-INT quantized 405B models on the Open LLM Leaderboard V2 compared to FP8, our approach reduces this to just 0.7 points.
 3 Benchmark Design and Setup
 3.1 Datasets and Benchmarks
-Our accuracy evaluation suite is designed to encompass a wide range of inference scenarios and use cases.
-It includes a diverse set of tasks to ensure a representative collection of topics relevant to practical deployments.
-Furthermore, the evaluation process is designed to be completely automated, as opposed to human-in-the-loop approaches, enabling streamlined experimentation.
-Overall, we group all of the benchmarks into three distinct categories.
-Academic benchmarks such as Open LLM Leaderboard V1 and V2 Beeching et al.
+We categorize benchmarks into three groups: academic, real-world, and text similarity analysis.
+1.
+Academic benchmarks, such as Open LLM Leaderboard V1 and V2 Beeching et al.
 (2023); Fourrier et al.
-(2024), are essential for evaluating model improvements.
-They focus on structured tasks like question-answering and reasoning, providing consistent and reproducible scoring.
-However, they often lack alignment with real-world scenarios where semantics, variability, and context-awareness are important.
-The Open LLM Leaderboard V1 consists of a diverse range of topics, including grade school math (GSM Cobbe et al.
-(2021)), world knowledge and reasoning (MMLU Hendrycks et al.
-(2020), ARC-Challenge Clark et al.
-(2018)), language understanding (Winogrande Sakaguchi et al.
-(2021), HellaSwag Zellers et al.
-(2019)), and truthfulness (TruthfulQA Lin et al.
-(2021)).
-The Open LLM Leaderboard V2 builds on this with topics such as expert knowledge and complex reasoning (MMLU-Pro Wang et al.
+(2024), provide structured evaluations for question-answering and reasoning tasks.
+While widely used for benchmarking, they lack alignment with real-world scenarios involving semantics, variability, and context-awareness.
+Leaderboard V1 includes tasks like GSM for grade school math Cobbe et al.
+(2021), MMLU and ARC-Challenge for world knowledge and reasoning Hendrycks et al.
+(2020); Clark et al.
+(2018), Winogrande and HellaSwag for language understanding Sakaguchi et al.
+(2021); Zellers et al.
+(2019), and TruthfulQA for factual correctness Lin et al.
+(2021).
+Leaderboard V2 extends this with expert knowledge benchmarks such as MMLU-Pro Wang et al.
 (2024), GPQA Rein et al.
-(2023), Big Bench Hard Suzgun et al.
-(2022)), multi-step reasoning (MuSR Sprague et al.
-(2024)), advanced math problems (MATH Level 5 Hendrycks et al.
+(2023), and Big Bench Hard Suzgun et al.
+(2022), as well as multi-step reasoning (MuSR Sprague et al.
+(2024)), advanced math (MATH Level 5 Hendrycks et al.
 (2021)), and instruction following (IFEval Zhou et al.
 (2023)).
-By evaluating with both Leaderboards, we cover a broad spectrum of topics.
-We employ log-likelihood (multiple-choice) and text-generation evaluations, allowing for stress testing of quantized LLMs under diverse conditions.
-Real-world benchmarks, unlike academic benchmarks, test models in scenarios that mimic human usage, such as instruction following, chat, and code generation.
-These benchmarks include Arena-Hard-Auto-v0.1 Li et al.
-(2024b); Chiang et al.
-(2024); Li et al.
-(2024c), HumanEval Chen et al.
-(2021), and HumanEval+ Liu et al.
-(2023a), which offer a broader range of tasks with higher variation but better reflect real-world environments.
-The LMSYS Chatbot Arena Chiang et al.
-(2024) has established itself as a leading benchmark for LLMs, assessing how models align with human preferences.
-Arena-Hard-Auto-v0.1 is an automated extension of this benchmark, where an LLM judges responses to 500 complex prompts on various topics.
-It has demonstrated a strong correlation with human evaluations, achieving a state-of-the-art 89% agreement with human preference rankings Li et al.
-(2024c).
-This enables a reliable, fast, and automated evaluation of the model’s chat capabilities without requiring a human.
-In addition to chat-based interactions, LLMs are widely deployed as coding assistants.
-To evaluate the performance of quantized models in code generation, we tested them on HumanEval and its more challenging variant, HumanEval+.
-These benchmarks measure a model’s ability to generate correct and functional code based on programming problems, with HumanEval+ introducing more rigorous evaluation protocol.
-Text similarity analysis evaluates how closely the generated outputs of quantized models align with those of their unquantized counterparts.
-While real-world benchmarks reflect practical usage scenarios for LLMs, their inherently open-ended format introduces notable variance in results.
-To address this, we extend our evaluation suite to include an analysis of output similarity under identical prompt requests.
-This allows us to get insights into how quantized models’ text generation differs from that of full-precision models.
-For this investigation, we employ metrics such as ROUGE Lin (2004), BERTScore Zhang et al.
-(2019), and Semantic Textual Similarity (STS) to assess both semantic and structural consistency, ensuring that the intended meaning and quality of generated text are preserved under quantization.
-ROUGE-1 measures word-level (unigram) overlap between outputs from quantized and unquantized models, while ROUGE-L captures structural similarity by focusing on the longest common subsequence.
-BERTScore evaluates token-level contextual similarity, computed using the RoBERTa-large model’s contextual embeddings.
-STS assesses overall semantic similarity at the sentence level, using the Sentence Transformers framework Reimers & Gurevych (2019) built on the 6-layer MiniLM architecture Wang et al.
-(2020).
-With this extensive evaluation framework, we ensured that deployment scenarios ranging from structured, research-driven tasks to open-ended, real-world applications were covered, providing a holistic view of the performance and capabilities of quantized LLMs.
-3.2 Models, Formats, and Quantization Algorithms
-We focus our evaluation on extensive testing of the recent Llama 3.1 series of models Dubey et al.
-(2024), which have gained significant traction in research and deployment contexts.
-To obtain a comprehensive understanding of trade-offs, we conduct all of our experiments on the instruction-tuned versions of all three available Llama 3.1 sizes (8B, 70B, and 405B), and for each of these, we focus on investigating three popular quantization formats with efficient inference support in vLLM: W8A8-FP, W8A8-INT, and W4A16-INT.
+By evaluating across both leaderboards, we capture a broad spectrum of reasoning and knowledge domains, using both log-likelihood and text-generation evaluations to stress-test quantized models.
 Table 1: Comparison of GPTQ and AWQ 4-bit weight quantization algorithms (W4A16-INT).
@@
 AWQ Lin et al. (2024a) 62.53 83.96 41.09 72.3 56.7 79.4 80.8
-W8A8-FP This scheme uses an 8-bit floating point format for weights and activations of all linear operators within transformer blocks.
-We compress models to this format using a simple round-to-nearest algorithm.
-Weight quantization follows a symmetric per-output-channel scheme, while activations are dynamically quantized per token.
-Due to the algorithm’s simplicity, this format requires no calibration data and is efficient even for massive models.
-W8A8-INT This scheme reduces the weights and activations of all linear operators within transformer blocks to 8-bit integer values.
-We apply symmetric per-output-channel GPTQ quantization for weight compression, and for activations, we use dynamic per-token quantization.
-We found that this scheme performs consistently well across 8B and 405B models, though it results in noticeable accuracy drops for the 70B model.
-To address this, we employ SmoothQuant, which redistributes the complexity of activation quantization onto the generally easier-to-quantize weights.
-For calibration data, we found that random tokens are adequate for the 8B scale; however, for larger models, we found that using the high-quality OpenPlatypus dataset Lee et al.
-(2023) is crucial for consistent accuracy recovery of competitive scores on academic benchmarks.
-W4A16-INT In this scheme, the weights of all linear operators within transformer blocks are quantized to 4-bit integers, while activations are retained at 16-bit precision.
-Weights are compressed using GPTQ quantization applied to groups of 128 consecutive elements with MSE-optimal clipping factors.
-Due to higher compression, using random tokens proved detrimental to accuracy; therefore, we employ calibration data sampled from the OpenPlatypus dataset.
+2.
+Real-world benchmarks evaluate models in practical scenarios such as instruction following, chat, long-context, and code generation.
+Arena-Hard-Auto-v0.1 (Li et al., 2024b; Chiang et al., 2024; Li et al., 2024c) automates LMSYS Chatbot Arena (Chiang et al., 2024) evaluations, using an LLM to judge responses to 500 complex prompts, achieving an 89% agreement with human rankings (Li et al., 2024c).
+This allows rapid and scalable assessment of chat capabilities without human intervention.
+For code generation, we evaluate models on HumanEval (Chen et al., 2021) and its extension HumanEval+ (Liu et al., 2023a), which test the ability to generate correct and functional code.
+Finally, we conduct long-context evaluations via the rigorous RULER benchmark (Hsieh et al., 2024) which consists of retrieval, multi-hop tracing, information aggregation, and question answering evaluations at sequence lengths from 4k to 128k.
+3.
+Our text similarity analysis benchmark assesses how closely quantized models’ outputs align with their full-precision counterparts.
+While real-world benchmarks reflect practical usage, their open-ended nature introduces variability, making direct accuracy comparisons challenging.
+To mitigate this, we analyze output similarity under identical prompts using ROUGE (Lin, 2004), BERTScore (Zhang et al., 2019), and Semantic Textual Similarity (STS) (Reimers and Gurevych, 2019).
+ROUGE-1 measures unigram overlap, while ROUGE-L captures structural similarity through the longest common subsequence.
+BERTScore computes token-level contextual similarity using RoBERTa-large embeddings, and STS assesses semantic alignment at the sentence level via Sentence Transformers built on MiniLM (Wang et al., 2020).
+3.2 Models, Formats, and Algorithms
+We evaluate using the highly-popular Llama 3.1 model series Dubey et al.
+(2024).
+To assess quantization trade-offs, we conduct experiments on the instruction-tuned versions of all available sizes (8B, 70B, and 405B).
+For each, we examine the three main formats with kernel support in vLLM: W8A8-FP, W8A8-INT, and W4A16-INT.
+W8A8-FP quantizes all linear operators in transformer blocks to an 8-bit floating-point format, using round-to-nearest quantization.
+Weights follow a symmetric per-output-channel scheme, while activations are dynamically quantized per token.
+This requires no calibration data and remains computationally efficient, even for large-scale models.
+W8A8-INT reduces weights and activations to 8-bit integers, applying symmetric per-output-channel GPTQ quantization for weights and dynamic per-token quantization for activations.
+While this scheme performs well for 8B and 405B models, it causes noticeable accuracy drops at 70B.
+To mitigate this, we apply SmoothQuant, shifting some activation complexity onto weights, which are easier to quantize.
+For calibration, random tokens suffice at 8B, but larger models require higher-quality calibration data, for which we use Lee et al.
+(2023).
+W4A16-INT quantizes weights to 4-bit integers while keeping activations at 16-bit precision.
+Weights are compressed using GPTQ with MSE-optimal clipping, applied in 128-element groups.
+Unlike higher-bit formats, random token calibration degrades accuracy, so we rely on OpenPlatypus data for calibration.
 INT4 Quantization Algorithms.
-As shown by prior work Dettmers et al.
-(2022); Lin et al.
-(2024a); Xiao et al.
-(2022), round-to-nearest performs poorly for integer weight quantization.
-Therefore, we focus on two very popular techniques with efficient inference support in vLLM: AWQ and GPTQ.
-We examine their performance on several benchmarks: Leaderboard V1 and V2, Arena-Hard, HumanEval, and MBPP.
-We report results in Table 1, and provide detailed per-task breakdown of scores in Appendix Tables 1213, and  8.
-We observe that, on average, on academic benchmarks, the methods are essentially tied.
-The advantage of AWQ, on a scale of 0–100, is only 0.23 and 0.35, respectively.
-However, we found that, on real-world benchmarks, GPTQ consistently outperforms AWQ by larger margins of 2.9 and 0.8, respectively.
-Consequently, we adopt GPTQ as our primary method for weight quantization.
-This conclusion is slightly at odds with prior work (Lin et al., 2024a; Huang et al., 2024) observing results that are either in favor of AWQ, or tied across a sub-sample of academic benchmarks.
-This is explained as follows: (1) contrary to this prior work, we use the version of GPTQ with MSE-optimal clipping factors (as opposed to min-max optimal), which has no overhead and performs consistently better; (2) we use different, higher-quality calibration data; (3) we evaluate across real-world benchmarks.
-4 Quantization versus Accuracy
-We begin our discussion by examining the accuracy of quantized models across Leaderboard V1 (Table 2), Leaderboard V2 (Table 3) and real-world benchmarks (Table 3).
-Given the density of the results, we examine them individually via average recoveries across higher-level benchmarks and discuss “outlier” observations.
+We focus on two inference-efficient techniques: AWQ and GPTQ, evaluating them on Leaderboard V1/V2, Arena-Hard, HumanEval, and MBPP.
+Results (Table 1) show near-identical performance on academic benchmarks, with AWQ leading by just 0.23 and 0.35 points on a 0–100 scale.
+However, GPTQ outperforms AWQ on real-world tasks by wider margins (2.9 and 0.8 points, respectively), leading us to adopt GPTQ as our primary INT4 method.
+This finding contrasts with prior studies (Lin et al., 2024a; Huang et al., 2024), which favored AWQ or found it tied on academic subsets.
+We attribute this to three key factors: (1) we use GPTQ with MSE-optimal clipping (the AWQ comparison used absmax); this has no overhead and yields consistently better results; (2) we use higher-quality calibration data than the C4 default; (3) we include real-world benchmarks, providing a broader evaluation scope.
 Table 2: Detailed per-task breakdown of accuracy on a subset of academic benchmarks (Open LLM Leaderboard V1) for quantized Llama-3.1-Instruct models across all three model sizes (8B, 70B, 405B).
@@
 W4A16-INT 199.98 86.78 87.2 87.7 95.3 96.3 88.3 87.4 65.3
-Table 3: Detailed per-task breakdown of accuracy on a subset of academic (Open LLM Leaderboard V2) and on real-world (Arena-Hard, HumanEval) benchmarks for quantized Llama-3.1-Instruct models across all three model sizes (8B, 70B, 405B).
-A higher score is better.
+Table 3: Detailed per-task breakdown of accuracy on a subset of academic (Open LLM Leaderboard V2) and on real-world (Arena-Hard, HumanEval, RULER) benchmarks for quantized Llama-3.1-Instruct models across all three model sizes (8B, 70B, 405B).
+Higher score is better.
+Long-context RULER evaluations at 405B are prohibitively expensive for our cluster.
 Academic Benchmarks (Open LLM Leaderboard V2) Real-World Benchmarks
-Recovery % Average Score IFEval 0-shot BBH 3-shot Math lvl 5 4-shot GPQA 0-shot MuSR 0-shot MMLU-Pro 5-shot Arena-Hard Win-Rate HumanEval pass@1 HumanEval+ pass@1
-8B BF16 100.00 27.62 77.86 30.09 15.68 3.68 7.61 30.77 25.8 67.3 60.7
-W8A8-FP 101.23 27.96 77.16 29.65 16.53 5.74 7.52 31.18 26.8 67.3 61.3
-W8A8-INT 101.52 28.04 77.95 30.96 15.51 5.38 7.61 30.85 27.2 67.1 60.0
-W4A16-INT 196.05 26.53 76.30 28.91 14.80 4.04 6.33 28.81 24.0 67.1 59.1
-70B BF16 100.00 41.66 86.41 55.79 26.07 15.40 18.16 48.12 57.0 79.7 74.8
-W8A8-FP 100.00 41.66 87.57 54.92 28.03 14.58 17.16 47.70 57.7 80.0 75.0
-W8A8-INT 197.29 40.53 86.56 55.17 23.94 13.61 16.82 47.09 57.0 78.7 74.0
-W4A16-INT 197.41 40.58 85.74 55.01 24.38 13.85 17.25 47.25 56.3 80.5 74.2
-405B BF16 100.00 48.73 87.69 67.02 38.96 19.50 19.48 59.74 67.4 86.8 80.1
-W8A8-FP 199.86 48.66 86.81 67.10 38.84 18.99 20.83 59.40 66.9 87.0 81.0
-W8A8-INT 198.36 47.93 86.99 66.73 35.82 20.38 19.24 58.43 64.6 86.9 80.4
-W4A16-INT 198.99 48.24 88.03 67.51 37.61 17.52 19.39 59.35 66.5 85.1 78.9
+Recovery % Average Score IFEval 0-shot BBH 3-shot Math lvl 5 4-shot GPQA 0-shot MuSR 0-shot MMLU-Pro 5-shot Arena-Hard Win-Rate HumanEval pass@1 HumanEval+ pass@1 RULER Score
+8B BF16 100.0 27.6 77.8 30.1 15.7 3.7 7.6 30.8 25.8 67.3 60.7 82.8
+W8A8-FP 101.2 27.9 77.2 29.6 16.5 5.7 7.5 31.2 26.8 67.3 61.3 82.8
+W8A8-INT 101.5 28.0 77.9 30.9 15.5 5.4 7.6 30.9 27.2 67.1 60.0 82.8
+W4A16-INT 196.1 26.5 76.3 28.9 14.8 4.1 6.3 28.8 24.0 67.1 59.1 81.1
+70B BF16 100.0 41.7 86.4 55.8 26.1 15.4 18.1 48.1 57.0 79.7 74.8 83.3
+W8A8-FP 100.0 41.7 87.6 54.9 28.0 14.6 17.2 47.7 57.7 80.0 75.0 83.0
+W8A8-INT 197.3 40.5 86.6 55.2 23.9 13.6 16.8 47.1 57.0 78.7 74.0 82.5
+W4A16-INT 197.4 40.6 85.7 55.0 24.4 13.8 17.2 47.2 56.3 80.5 74.2 82.2
+405B BF16 100.0 48.7 87.7 67.0 38.9 19.5 19.5 59.7 67.4 86.8 80.1 -
+W8A8-FP 199.9 48.7 86.8 67.1 38.8 18.9 20.8 59.4 66.9 87.0 81.0 -
+W8A8-INT 198.3 47.9 86.9 66.7 35.8 20.4 19.2 58.4 64.6 86.9 80.4 -
+W4A16-INT 198.9 48.2 88.0 67.5 37.6 17.5 19.4 59.3 66.5 85.1 78.9 -
+4 Quantization Impact on Accuracy
+We begin our discussion of the results by examining the accuracy of quantized models across Leaderboard V1 (Table 2), Leaderboard V2 (Table 3) and real-world benchmarks (Table 3).
+Given the density of the results, we discuss them individually via average recoveries across higher-level benchmarks and discuss “outlier” observations.
 4.1 Academic Benchmarks
-Academic benchmarks are the standard approach for evaluating LLM reasoning capabilities.
-They provide structured tasks with well-defined evaluation criteria.
-Our evaluations focus on the Open LLM Leaderboard V1 and V2 benchmarks.
-Testing on both allowed us to prevent overfitting to V1, where we optimized our quantization hyperparameters.
-The Open LLM Leaderboard V1 was evaluated by following Meta’s prompts for Llama-3.1 models to align with the expected baseline scores.
-This approach introduces two notable differences from the standard protocols: first, MMLU and ARC-Challenge are evaluated as text-generation tasks where models generate the correct answer as text, as opposed to the default log-likelihood-based protocol Gao et al.
-(2021).
-Second, GSM8k is evaluated using a chain-of-thought rather than the default few-shot format.
-As shown in Figure 1 and Table 2, all quantization schemes, regardless of model size, recover around 99% of the average score achieved by the unquantized BF16 baseline.
-Appendix Table 10 provides per-task accuracy recoveries, indicating that the lowest recovery is 96.88% for the W4A16-INT 8B model on the TruthfulQA task.
-For larger models, the lowest recovery also occurs on TruthfulQA, reaching approximately 98.5%.
-We observe that the average accuracy recovery across 8-bit quantization schemes is approximately 99.75%, indicating a negligible impact on model accuracy.
-Similarly, the W4A16-INT scheme achieves a highly competitive average accuracy recovery of 99.36%.
-Figure 1: Average score of quantized Llama-3.1-Instruct models on the Open LLM Leaderboard V1 benchmark.
-The Open LLM Leaderboard V2 was developed as scores on Leaderboard V1 reached a plateau.
-This benchmark incorporates more challenging tasks designed to assess advanced reasoning capabilities.
-Unlike V1, Leaderboard V2 normalizes scores by subtracting the random accuracy baseline and rescaling the adjusted score to a 0-100 range.
-This normalization provides a more robust method for calculating the aggregate average score by equally weighting each task, independent of its random prediction baseline.
-As illustrated in Figure 2 and Table 3, quantized models recover close to 99% of the baseline’s average score, with all models maintaining at least 96% recovery.
-However, the increased difficulty of these tasks, especially for smaller models, resulted in higher variance for benchmarks like GPQA and MuSR, where scores approached the random guessing threshold even for the full-precision baseline.
-This led to more variability and a lack of a clear signal for accuracy recovery, as seen in Appendix Table 11.
-Focusing on tasks where the full-precision model scores above 40%, representing cases where the model performs significantly better than random guessing, we observe that the lowest per-task accuracy recovery for 8-bit FP quantized models is 98.44% at the 70B scale on the Big Bench Hard (BBH) dataset.
-For 8-bit INT models, the lowest recovery is 97.8% at the 405B scale on the MMLU-Pro task.
-Notably, in this benchmarking setup, W4A16-INT models exhibit higher accuracy recovery than W8A8-INT models, with the lowest recovery being 98% for the 8B model on the IFEval task.
-Figure 2: Average score of quantized Llama-3.1-Instruct models on the Open LLM Leaderboard V2 benchmark.
-4.2 Real-World Benchmarks
-While academic benchmarks provide structured evaluations, real-world open-ended benchmarks better represent how models perform in dynamic environments.
-These benchmarks test models on varied prompts with longer generations and multiple potential solutions, focusing on responses’ correctness and semantic quality.
-Our evaluations targeted three key real-world benchmarks: Arena-Hard-Auto-v0.1, HumanEval, and HumanEval+, which measure performance in chat, instruction-following, and code generation.
-Table 3 summarizes results on these benchmarks.
-Figure 3: Average score of quantized Llama-3.1-Instruct models on the Arena-Hard-Auto-v0.1.
-Figure 3 focuses on the Arena-Hard-Auto-v0.1 benchmark, averaging results from two evaluation runs per model and quantization scheme.
-The results illustrate that the response quality of quantized models remains highly competitive with their unquantized counterparts.
-As shown in the detailed results in Appendix Table 7, the 95% confidence intervals overlap for all model sizes and quantization schemes, highlighting the minimal impact of quantization on accuracy.
-Figures 4 and 5 present the pass@1 scores obtained using the EvalPlus library Liu et al.
-(2023a).
-As illustrated, quantized models demonstrate exceptional performance on both HumanEval and HumanEval+, with 8-bit models achieving 99.9% accuracy recovery and 4-bit models recovering 98.9%.
-These results highlight that quantized models maintain high performance in both simpler and more complex coding tasks, proving their reliability for real-world applications.
-In Appendix A.1, we report results for pass@10 metric, which show similar trends to the pass@1 metric.
-Figure 4: HumanEval pass@1 scores for quantized Llama-3.1-Instruct models.
-Figure 5: HumanEval+ pass@1 scores for quantized Llama-3.1-Instruct models.
+Our first analysis focuses on Open LLM Leaderboard V1 and V2, ensuring generalization by optimizing quantization hyperparameters on V1 while validating results on V2.
+The Open LLM Leaderboard V1 follows Meta’s prompt guidelines for Llama-3.1 models to maintain alignment with baseline scores.
+This introduces two key differences from standard evaluation protocols: MMLU and ARC-Challenge are assessed as text-generation tasks rather than log-likelihood-based evaluations Gao et al.
+(2021), and GSM8k is tested using chain-of-thought prompting instead of a few-shot approach.
+Table 2 shows that all quantization schemes, across model sizes, recover approximately 99% of the unquantized BF16 baseline.
+The lowest task recovery occurs on TruthfulQA, reaching 96.88% for W4A16-INT at 8B and $\sim$98.5for larger models (see Appendix Table 9).
+On average, 8-bit quantization achieves 99.75% recovery, while W4A16-INT reaches a competitive 99.36%.
+The Open LLM Leaderboard V2 incorporates more challenging tasks to assess advanced reasoning.
+Unlike V1, V2 normalizes scores by subtracting the random baseline and rescaling to a 0-100 range, ensuring equal weighting across tasks regardless of inherent difficulty.
+Table 3 shows that quantized models maintain 99% of the baseline’s average score, with all models recovering at least 96%.
+However, due to the increased difficulty, smaller models exhibit higher variance, particularly on GPQA and MuSR, where full-precision models already approach random guessing thresholds, reducing the reliability of accuracy recovery signals (Appendix Table10).
+Focusing on tasks where the full-precision model scores above 40%, ensuring a meaningful performance baseline, we observe the lowest per-task recovery for 8-bit FP quantization at 98.44% on BBH (70B) and for 8-bit INT at 97.8% on MMLU-Pro (405B).
+Notably, W4A16-INT models demonstrate superior recovery over W8A8-INT, with a minimum accuracy retention of 98% for the 8B model on IFEval.
+This suggests that, for INT, quantizing activations is harder than quantizing weights.
 Figure 1: Text similarity metrics comparing the outputs of quantized Llama-3.1-Instruct models to full-precision baselines. We refer to W8A8-FP as FP8, W8A8-INT as INT8, and W4A16-INT as INT4.
+4.2 Accuracy on Real-World Benchmarks
+While academic benchmarks offer structured evaluations, real-world benchmarks better capture model performance in dynamic environments.
+These evaluations involve diverse prompts, longer generations, and multiple valid responses, emphasizing correctness and semantic quality.
+We assess four key benchmarks: Arena-Hard-Auto-v0.1 (measuring chat and instruction-following performance, averaging two runs per model and quantization scheme), HumanEval, and HumanEval+ (measuring code generation quality and reporting pass@1 scores using the EvalPlus library Liu et al.
+(2023a)), and RULER (evaluating long-context abilities).
+Table 3 summarizes the results.
+On Arena-Hard-Auto-v0.1, quantized models exhibit competitive response quality, with overlapping 95% confidence intervals across all configurations (Appendix Table 6).
+In coding evaluations, quantized models also maintain strong performance, with 8-bit achieving 99.9% recovery and 4-bit recovering 98.9%, demonstrating their robustness across simple and complex coding tasks.
+Similarly, for the long-context RULER benchmark, quantized models achieve average score recovery of $\geq$ 98% across all formats.
+See Appendix A.1 for additional results.
 4.3 Text Similarity Investigation
-Following our investigations on various structured and open-ended tasks, we focus on examining the similarity of generated text between quantized models and their unquantized counterparts.
-As described before, we employ four metrics: ROUGE-1, ROUGE-L, BERTScore, and Semantic Textual Similarity (STS).
-All of these metrics were computed across responses generated to the Arena-Hard-Auto-v0.1 prompts with a greedy sampling technique, allowing us to analyze how well-quantized models preserve the meaning and structure of outputs compared to full-precision models on challenging prompts while ensuring full reproducibility.
-For visualization purposes, we normalize STS score to 0-1 range.
-Results in Figure 6 show that large quantized models (70B and 405B) maintain a high degree of similarity to their full-precision counterparts, with an average ROUGE-1 score of 0.7 and ROUGE-L score of 0.56, indicating strong preservation of word choice and structure.
-Additionally, an average BERTScore of 0.93 and STS score of 0.96 confirm that the overall meaning is consistent, despite minor token variations from quantization.
-While the 8B models exhibit slightly more variability in word choice, reflected in lower ROUGE-1 and ROUGE-L scores of 0.62 and 0.46, respectively, they still preserve core semantic meaning, as evidenced by the high BERTScore of 0.92 and STS score of 0.95.
-These findings demonstrate that quantized models produce high-quality outputs across all analyzed model sizes and schemes.
-5 Quantization and Inference Performance
-LLM inference occurs in two main stages: prefill, where all tokens in the input prompt are processed simultaneously, and decode, where new tokens are generated one at a time.
-For most input prompts, prefill is compute-bound, while decode tends to be memory-bound.
-Weight quantization primarily enhances performance in the decode stage by reducing memory load times.
-Conversely, quantizing both weights and activations allows matrix multiplications to be executed at lower bitwidth, boosting computational speed and reducing bottlenecks during prefill.
-Therefore, the optimal quantization scheme for inference depends on the distribution of prefill versus decode tokens.
-In addition, quantization also offers end-to-end performance benefits by enabling more simultaneous queries, which improves efficiency, and by facilitating the use of lower-cost GPUs for tasks with tighter memory constraints.
-These factors interact during real-world LLM deployment, making the choice of “optimal” quantization format complex.
-In this study, we benchmarked three quantization schemes (W8A8-FP, W8A8-INT, and W4A16-INT) across three GPU types (A6000, A100, H100) in six latency-sensitive use cases and two throughput use cases, as listed in Table 4.
-We considered both synchronous deployment and asynchronous deployment with latency constraints for latency-sensitive use cases, while throughput use cases were evaluated with asynchronous deployment only.
-We release the inference-benchmarking data as a HuggingFace dataset.
-Furthermore, to examine the cost-efficiency of deployment in different hardware, we use Lambda Labs’ on-demand GPU pricing Lambda Labs (2024), which are representative of industry standards.
-The prices are reported in Table 9.
-Table 4: Use cases for benchmarking the performance of quantized models.
-For latency-sensitive use cases, we impose the following constraints: TTFT $<$ 1s and ITL $<$ 0.1s (TTFT = time to first token, ITL = inter-token latency).
-Latency-Sensitive Use Cases
-Use Case Prefill Tokens Decode Tokens
-Code Completion 1256 1024
-Docstring Generation 1768 1128
-Code Fixing 1024 1024
-RAG 1024 1128
-Instruction Following 1256 1128
-Multi-Turn Chat 1512 1256
-Throughput Use Cases
-Small Summarization 1024 1128
-Large Summarization 4096 1512
+Next, we analyze the similarity of generated text between quantized and full-precision models.
+Using Arena-Hard-Auto-v0.1 prompts and greedy sampling for full reproducibility, we compute ROUGE-1, ROUGE-L, BERTScore, and Semantic Textual Similarity (STS) normalized to a 0-1 range.
+As shown in Figure 1, large quantized models (70B and 405B) closely match their full-precision counterparts, achieving an average ROUGE-1 of 0.7 and ROUGE-L of 0.56, indicating strong word and structural preservation.
+BERTScore (0.93) and STS (0.96) further confirm semantic consistency despite minor token variations.
+While 8B models exhibit slightly higher variability, with ROUGE-1 and ROUGE-L dropping to 0.62 and 0.46, they still maintain strong semantic fidelity, as reflected in their BERTScore (0.92) and STS (0.95).
+These results demonstrate that quantized models generate high-quality outputs across all sizes and schemes.
 Table 4: Synchronous inference performance comparison across model sizes and GPU configurations.
@@
 We refer to W8A8-FP as FP8, W8A8-INT as INT8, and W4A16-INT as INT4.
-Model GPU # Format CR Code Completion Docstring Generation Code Fixing RAG Instruction Following Multi-Turn Chat
-Lat. Q/\$ Lat. Q/\$ Lat. Q/\$ Lat. Q/\$ Lat. Q/\$ Lat. Q/\$
-8B A6000 1 BF16 7.7 584 3.2 1,403 7.2 629 3.3 1,345 3.3 1,383 5.7 792
-1 INT8 1.34 6.0 746 2.2 2,011 8.1 554 2.3 1,955 1.9 2,327 4.3 1,058
-1 INT4 2.09 3.7 1,201 1.6 2,896 4.8 933 1.6 2,795 1.2 3,623 2.6 1,752
-70B A6000 4 BF16 8.7 129 8.7 129 12.8 88 8.0 141 6.5 172 13.6 83
-2 INT8 1.84 14.0 160 8.3 272 13.4 167 8.3 270 7.4 303 12.7 177
-2 INT4 3.09 8.2 276 5.2 429 7.6 296 5.8 390 4.5 502 6.7 334
-A100 2 BF16 7.9 128 6.3 161 11.2 90 6.1 164 5.3 190 9.7 104
-1 INT8 1.73 11.5 175 6.6 306 11.9 170 7.0 290 6.4 312 10.8 186
-1 INT4 2.75 6.3 319 4.7 428 6.8 297 5.1 396 3.9 510 6.6 306
-H100 2 BF16 4.9 115 3.8 147 7.0 81 3.9 144 3.4 165 5.7 99
-1 FP8 1.89 5.2 212 4.0 272 7.0 156 4.0 272 3.5 317 5.9 185
-1 INT4 2.51 4.5 241 3.1 358 4.5 244 3.4 325 2.7 411 4.0 271
-405B A100 16 BF16 22.8 6 11.1 11 26.6 5 12.2 10 10.1 12 20.5 6
-8 INT8 3.35 14.0 18 6.7 37 15.7 16 6.8 37 6.4 39 12.0 21
-4 INT4 7.28 13.1 38 6.9 72 11.4 44 7.3 69 6.2 81 10.7 47
-H100 16 BF16 15.0 5 6.7 11 17.7 4 6.7 11 6.5 12 12.2 6
-8 FP8 3.45 7.8 19 4.2 36 8.8 17 4.2 36 4.1 37 7.5 20
-4 INT4 6.75 7.8 37 4.6 63 7.5 39 4.8 61 4.0 72 7.0 42
+Size GPU # Format CR Code Completion Docstring Generation Code Fixing RAG Instruction Following Multi-Turn Chat Summarization
+Lat. Q/\$ Lat. Q/\$ Lat. Q/\$ Lat. Q/\$ Lat. Q/\$ Lat. Q/\$  Lat. Q/\$
+8B A6000 1 BF16 24.5 183 3.2 1,395 25.0 180 3.3 1,374 3.1 1,445 6.2 723 13.4 335
+1 INT8 1.54 15.9 284 2.1 2,157 16.3 276 2.1 2,139 2.0 2,249 4.0 1,120 8.9 506
+1 INT4 2.39 9.7 462 1.4 3,290 10.1 445 1.4 3,136 1.3 3,543 2.5 1,787 6.1 736
+70B A6000 4 BF16 61.7 18 6.6 170 62.6 18 8.1 138 8.0 141 15.8 71 32.6 35
+2 INT8 1.94 63.4 35 7.1 317 63.8 35 8.4 267 8.0 280 16.2 139 34.0 66
+2 INT4 2.96 39.2 57 5.0 453 40.4 56 5.8 390 5.1 440 10.2 221 23.5 96
+A100 2 BF16 50.7 20 2.9 343 51.2 20 6.8 148 6.4 156 12.9 78 27.3 37
+1 INT8 1.81 54.3 37 4.0 500 54.8 37 7.2 279 6.9 291 13.8 146 29.3 69
+1 INT4 2.67 35.0 57 2.8 718 35.8 56 5.2 390 4.6 439 9.2 220 21.0 96
+H100 2 BF16 31.3 18 4.0 139 31.5 18 4.1 138 4.0 142 7.9 71 16.4 34
+1 FP8 1.84 32.8 33 4.3 256 33.1 33 4.3 254 4.2 262 8.3 132 17.4 63
+1 INT4 2.11 28.6 38 3.8 289 28.2 39 3.8 287 3.7 299 7.1 153 15.3 72
+405B A100 16 BF16 81.9 2 10.8 12 81.2 2 11.2 11 10.6 12 20.9 6 44.1 3
+8 INT8 3.27 50.1 5 6.6 38 50.5 5 6.8 37 6.4 39 12.8 20 26.9 9
+4 INT4 6.38 48.9 10 7.0 71 49.5 10 7.3 68 6.4 79 12.7 39 29.4 17
+H100 16 BF16 50.6 1 6.5 12 50.3 1 6.6 11 6.4 12 13.0 6 26.5 3
+8 FP8 3.17 31.7 5 4.2 36 31.9 5 4.2 36 4.1 37 8.0 19 16.7 9
+4 INT4 5.15 37.5 8 5.0 58 37.8 8 5.1 57 4.8 60 9.2 32 20.4 14
 CR: Cost Reduction factor compared to BF16 baseline. Higher is better.
 Lat.: Latency in seconds (lower is better). Q/\$: Queries per USD (higher is better).
-5.1 Synchronous Deployment
-Latency-sensitive use cases are sometimes deployed in synchronous mode, i.e., only one query is processed at a time.
-Synchronous deployment leads to the lowest latency since different requests do not compete for computation, and it tends to be dominated by decode time.
-Table 5 compares the inference performance for synchronous deployment across different model sizes, GPU types, quantization schemes, and use cases.
-For each configuration, we highlight the number of GPUs that are most cost-effective.
-The results reveal that W4A16-INT quantization consistently delivers the highest performance gains for synchronous deployment across all model sizes, GPU types, and latency-sensitive use cases.
-For both the 8B and 70B model sizes, W4A16-INT achieves significant 2-3x cost reduction per query and 1.5-2x latency improvements compared to the baseline full-precision BF16 format.
-For the 405B model, W4A16-INT is exceptionally effective, offering 6–7x reductions in cost per query and enabling inference on a smaller number of GPUs.
-Specifically, deploying the 405B model with W4A16-INT on 4x A100 or H100 GPUs meets performance thresholds that previously required 16 GPUs in BF16 format, reducing inter-GPU communication and lowering latency.
-As seen in the previous section, the accuracy differences are minor.
-5.2 Asynchronous Deployment
-Processing multiple queries simultaneously generally results in better computational efficiency than a single query at a time.
-vLLM automatically coordinates multiple asynchronous queries, balancing computation between prefill and decode stages of parallel requests.
-Asynchronous deployment usually results in increased latency compared to synchronous deployment, but the processing time is amortized across multiple queries, leading to significant gains in average queries per second.
-For latency-sensitive use cases, we report the maximum inference throughput (queries per second) under the constraints listed in Table 4, whereas for “throughput” use cases, we consider the absolute maximum inference throughput.
-Table 6 shows the inference performance for the asynchronous deployment of both latency-sensitive and throughput use cases.
-The hardware configurations were optimized to match the number of GPUs yielding the best performance for full-precision BF16 deployment, and we assume optimal orchestration for the quantized models.
 Table 5: Asynchronous inference performance evaluation across model sizes and hardware configurations.
@@
 We refer to W8A8-FP as FP8, W8A8-INT as INT8, and W4A16-INT as INT4.
-Mdl. HW Fmt. Speedup Code Compl. Doc. Gen. Code Fixing RAG Inst. Following Multi-Turn Chat Sum. (S) Sum. (L)
-QPS Q/\$ QPS Q/\$ QPS Q/\$ QPS Q/\$ QPS Q/\$ QPS Q/\$ QPS Q/\$ QPS Q/\$
-8B 1×A6000 BF16 4.5 20.2k 5.5 24.8k 2.2 10.1k 4.2 18.9k 10.3 46.5k 5.3 23.7k 4.3 19.5k 1.0 4.3k
-INT8 1.16 4.5 20.3k 7.0 31.4k 1.9 8.7k 5.2 23.3k 16.6 74.5k 5.1 22.9k 5.8 26.3k 1.1 5.1k
-INT4 1.56 8.0 35.8k 9.5 42.8k 3.9 17.7k 7.4 33.4k 17.8 80.2k 8.7 39.2k 4.0 17.9k 0.9 3.9k
-70B 4×A6000 BF16 0.7 794 0.4 491 0.4 478 0.6 631 1.3 1.4k 0.5 537 1.4 1.6k 0.3 312
-INT8 3.10 2.4 2.7k 1.5 1.6k 1.2 1.4k 1.4 1.5k 3.6 4.1k 1.8 2.0k 2.5 2.8k 0.4 504
-INT4 3.17 3.1 3.5k 1.6 1.8k 0.8 947 0.7 814 3.8 4.3k 2.2 2.5k 3.1 3.5k 0.3 363
-4×A100 BF16 5.5 2.8k 3.1 1.6k 2.3 1.2k 2.6 1.3k 6.2 3.1k 3.6 1.8k 4.2 2.1k 1.0 514
-INT8 1.75 7.6 3.8k 5.7 2.9k 4.1 2.0k 4.7 2.4k 11.6 5.8k 6.7 3.4k 6.9 3.5k 1.7 844
-INT4 1.33 7.7 3.9k 4.0 2.0k 2.7 1.4k 3.0 1.5k 9.5 4.8k 5.2 2.6k 7.7 3.9k 0.9 443
-4×H100 BF16 11.7 3.4k 7.6 2.2k 5.4 1.6k 6.4 1.9k 13.3 3.9k 7.9 2.3k 8.8 2.6k 2.2 630
-FP8 1.67 19.4 5.6k 13.0 3.8k 8.9 2.6k 10.5 3.1k 22.9 6.7k 12.7 3.7k 19.4 5.6k 4.3 1.3k
-INT4 1.10 13.5 3.9k 7.9 2.3k 5.4 1.6k 5.9 1.7k 17.1 5.0k 9.5 2.8k 13.5 4.0k 1.6 466
-405B 16×A100 BF16 0.4 48 0.4 47 0.2 25 0.3 38 0.8 96 0.3 42 2.2 276 0.4 56
-INT8 8.09 3.7 461 2.1 269 1.8 228 2.1 270 6.0 755 3.1 390 3.7 470 0.9 119
-INT4 7.12 3.7 464 2.2 280 1.5 190 1.6 202 5.1 639 2.5 320 2.8 353 0.6 70
-16×H100 BF16 1.8 137 1.4 109 0.8 63 1.2 92 2.7 204 1.4 107 4.0 302 0.8 62
-FP8 5.61 8.3 627 8.2 618 4.9 366 7.1 533 15.1 1.1k 8.8 663 6.2 464 1.8 133
-INT4 3.53 7.2 542 4.4 330 3.1 235 3.4 257 10.1 757 5.6 423 4.6 344 1.0 77
-QPS: Queries per second (higher is better). Q/\$: Queries per USD (higher is better). Sum. (S/L): Small/Large Summarization.
+Size HW Format Speedup Code Compl. Doc. Gen. Code Fixing RAG Inst. Following Multi-Turn Chat Summarization
+QPS Q/\$ QPS Q/\$ QPS Q/\$ QPS Q/\$ QPS Q/\$ QPS Q/\$  QPS Q/\$
+8B 1×A6000 BF16 1.5 6.8k 5.6 25.1k 1.1 4.8k 4.4 19.9k 11.8 53.0k 5.3 24.0k 0.7 3.2k
+INT8 1.38 2.2 9.8k 7.7 34.6k 1.4 6.4k 6.1 27.6k 16.5 74.5k 7.2 32.3k 1.0 4.4k
+INT4 1.08 2.2 9.8k 5.3 24.0k 1.3 6.0k 4.1 18.6k 11.2 50.5k 5.4 24.3k 0.7 3.1k
+70B 4×A6000 BF16 0.4 0.4k 1.4 1.6k 0.3 0.3k 1.4 1.6k 3.3 3.8k 1.5 1.7k 0.2 0.3k
+INT8 1.91 0.7 0.8k 3.9 4.4k 0.5 0.6k 2.8 3.1k 6.9 7.7k 2.2 2.5k 0.3 0.4k
+INT4 1.92 1.2 1.4k 2.7 3.1k 0.7 0.8k 1.9 2.1k 5.2 5.9k 2.6 3.0k 0.3 0.3k
+4×A100 BF16 1.4 0.7k 6.9 3.5k 1.0 0.5k 3.3 1.6k 8.7 4.4k 4.3 2.2k 0.7 0.4k
+INT8 1.87 2.4 1.2k 15.9 8.0k 1.8 0.9k 6.1 3.1k 16.5 8.3k 8.0 4.0k 1.2 0.6k
+INT4 1.64 2.3 1.2k 22.8 11.5k 1.4 0.7k 4.3 2.2k 11.9 6.0k 5.8 2.9k 0.8 0.4k
+4×H100 BF16 3.5 1.0k 10.0 2.9k 2.6 0.7k 8.0 2.3k 20.3 5.9k 9.9 2.9k 1.7 0.5k
+FP8 1.77 6.9 2.0k 17.8 5.2k 4.0 1.2k 14.3 4.2k 38.3 11.1k 18.4 5.4k 2.6 0.8k
+INT4 1.55 5.9 1.7k 16.4 4.8k 3.1 0.9k 13.0 3.8k 35.8 10.4k 16.1 4.7k 2.2 0.6k
+405B 16×A100 BF16 0.8 59 2.5 187 0.3 20 2.1 156 4.6 347 2.1 158 0.3 22
+INT8 2.53 1.3 98 4.8 358 1.1 79 3.8 282 10.1 760 4.9 366 0.8 63
+INT4 2.21 1.9 144 3.6 271 1.2 93 2.8 211 8.2 616 4.0 304 0.6 43
+16×H100 BF16 0.7 52 6.1 456 0.6 44 4.8 363 8.5 638 5.3 398 0.6 46
+FP8 3.04 4.4 329 9.6 725 2.7 200 7.6 571 20.7 1561 10.4 780 1.7 125
+INT4 3.09 4.0 304 11.1 833 2.5 192 8.7 652 24.7 1856 11.6 872 1.6 122
+QPS: Queries per second (higher is better). Q/\$: Queries per USD (higher is better).
 Numbers denoted with k represent thousands (e.g., 20.3k = 20,300).
-The data indicates that W8A8-INT and W8A8-FP are the most effective quantization schemes for the asynchronous deployment of mid- and large-size models on high-end GPUs.
-As expected, W8A8-FP performs best on H100 GPUs, while W8A8-INT is more suited for efficient inference on A100 GPUs.
-In contrast, W4A16-INT is the best option for cost efficiency on A6000 GPUs.
-It outperforms other quantization schemes, especially for smaller models and latency-sensitive tasks.
-It suggests that W4A16-INT is a viable and cost-effective option for less memory-intensive applications on mid-tier hardware.
-When evaluating cost efficiency across different GPU classes for mid- and large-size models, W8A8-FP on H100 GPUs consistently stands out as the best option across all use cases.
-Overall, the results underscore no universally optimal quantization scheme for asynchronous deployment.
-The best choice depends on model size, GPU type, and whether the task prioritizes latency or throughput.
-Nonetheless, quantization leads to significant cost reduction per query regardless of the scenario, achieving 1.5-3x cost reduction for 8B and 70B and 5-8x for 405B models.
+5 Quantized Inference Performance
+LLM inference consists of two main stages: prefill, where all input tokens are processed simultaneously, and decode, where tokens are generated sequentially.
+Prefill is typically compute-bound, while decode is memory-bound.
+Weight quantization primarily accelerates decode by reducing memory movement, whereas weight-and-activation quantization improves computational efficiency in prefill.
+Thus, the optimal choice for quantization scheme depends on the ratio of prefill to decode tokens.
+Beyond direct speedups, quantization also enhances end-to-end performance by increasing the number of simultaneous queries, improving efficiency, and enabling lower-cost GPU usage for memory-constrained tasks.
+Thus, real-world deployment involves complex trade-offs.
+To assess these trade-offs, we benchmarked W8A8-FP, W8A8-INT, and W4A16-INT across three GPU types (A6000, A100, H100) in seven use cases.
+Tasks like code completion and instruction following involve short prefill phases (256 tokens) and varying decode lengths (1024 and 128 tokens, respectively).
+More complex tasks like summarization require significantly longer prefill (4096 tokens) with a moderate decode length (512 tokens).
+Multi-turn chat and RAG involve moderate prefill lengths (512 and 1024 tokens) with shorter decode phases (256 and 128 tokens).
+Finally, docstring generation (768 prefill, 128 decode) and code fixing (1024 prefill, 1024 decode) reflect intermediate token requirements.
+For latency-sensitive applications, we compare both synchronous and asynchronous deployment under latency constraints, while throughput-driven cases are evaluated in asynchronous mode.
+To assess cost efficiency across hardware setups, we use Lambda Labs’ on-demand GPU pricing Lambda Labs (2024), shown in Table 8, which is standard.
+5.1 Synchronous Deployment
+Latency-sensitive applications are sometimes deployed in synchronous mode, where a single query is processed at a time.
+This approach minimizes latency by avoiding resource contention, making inference largely decode-bound.
+Table 4 compares inference performance across model sizes, GPU types, quantization schemes, and use cases, highlighting the most cost-effective GPU configurations.
+The results show that W4A16-INT consistently achieves the highest performance gains across all models and hardware setups.
+For 8B and 70B models, W4A16-INT reduces cost per query by 2–3× and improves latency by 1.5–2.5× compared to the full-precision BF16 baseline.
+The impact is even more pronounced at 405B, where W4A16-INT achieves 5–7× cost reductions and enables inference with fewer GPUs.
+Notably, deploying the 405B model on 4× A100 or H100 GPUs with W4A16-INT meets performance thresholds that previously required 16 GPUs in BF16, reducing inter-GPU communication and latency.
+Given the minor accuracy trade-offs observed in the previous section, this makes W4A16-INT highly effective for synchronous deployment.
+5.2 Asynchronous Deployment
+Processing multiple queries concurrently improves computational efficiency compared to single-query execution.
+vLLM automatically manages asynchronous requests, balancing computation between prefill and decode stages.
+While asynchronous deployment increases per-query latency relative to synchronous execution, it amortizes computation across multiple requests, significantly boosting overall throughput, measured in queries per second (QPS).
+Table 5 reports the maximum achievable throughput and cost efficiency (queries per dollar) across different quantization formats, model sizes, and hardware configurations.
+The setups were optimized for peak BF16 performance and kept consistent when evaluating quantized models.
+Results show that W8A8-INT and W8A8-FP yield the highest throughput, though W4A16-INT remains competitive and can outperform W8A8 in some scenarios.
+Many real-world applications impose latency constraints on asynchronous deployment.
+Figures 2 and  3 illustrate trade-offs between latency and throughput for two example tasks: docstring generation and code fixing.
+W4A16-INT is more efficient at lower latencies, making it ideal for applications requiring rapid response times.
+In contrast, W8A8 formats maximize throughput at the cost of higher latency, making them better suited for batch processing.
+The point where W8A8 overtakes W4A16 depends on factors such as model size, hardware, and task requirements.
+Figure 2: Latency-throughput example for docstring generation use-case.
+W4A16 is more efficient at low latency (lower throughput), whereas W8A8 becomes more efficient at high latency (high throughput).
+Figure 3: Latency-throughput example for code fixing use-case. W4A16 is more efficient at low latency (lower throughput), whereas W8A8 becomes more efficient at high latency (high throughput).
+Figure 4: Accuracy recovery trends across academic benchmarks highlight the challenges of integer activation quantization, particularly at larger model sizes.
 6 Conclusion
-We provided a broad and in-depth study of accuracy-vs-performance-vs-cost trade-offs for quantized LLMs across various deployment environments.
-Specifically, we covered all quantization formats with efficient computational support and a range of quantization algorithms, deployment use cases, and GPUs.
+We provided a broad, in-depth study of accuracy-vs-performance-vs-cost trade-offs for quantized LLMs across various deployment environments, covering all quantization formats with efficient support, and a range of quantization algorithms, deployment use cases, and GPUs.
 In Figure 4 we summarize our findings in terms of accuracy recovery per quantization format, using carefully-tuned state-of-the-art quantization techniques.
-Figure 7: Accuracy recovery trends across academic benchmarks highlight the challenges of integer activation quantization, particularly at larger model sizes.
-Broadly speaking, our findings show that, with a judicious choice of algorithm and parametrization, these formats can offer higher accuracy than previously thought, significantly improve inference performance, and reduce costs.
+Broadly, our findings show that, with a judicious choice of algorithm and parametrization, these formats can offer higher accuracy than previously thought, significantly improve inference performance, and reduce costs.
 At the same time, we have also shown that the optimal choice of format can be task and algorithm specific, providing guidelines for this choice.
-Acknowledgements
-We would like to express our sincere gratitude to the Neural Magic team, the vLLM community, and EleutherAI for developing the libraries that enabled this research.
-Specifically, we used the LLM Compressor to produce most of our quantized models, vLLM as an inference engine to accelerate evaluations, and the LM Evaluation Harness to run all of our academic benchmarks.
+Limitations
+While our study provides a comprehensive evaluation of quantization effects on model accuracy and inference performance, several limitations remain.
+We have primarily focused on weight and activation quantization, leaving open questions about the impact of compressing other model components such as the KV-cache, input embeddings, and language modeling head.
+Further investigation is needed to assess how these additional compression techniques influence both accuracy and computational efficiency.
+Additionally, our analysis does not fully explore the effects of quantization across specialized use cases, such as multi-lingual tasks, where accuracy degradation could vary significantly depending on the language distribution and underlying model architecture.
+Future work should extend these evaluations to provide a more holistic understanding of quantization trade-offs in diverse deployment scenarios.
 Appendix A Additional Results
 A.1 Real-World Benchmarks
-In Tables 8 and 9 we report pass@10 scores for all models on HumanEval and HumanEval+ benchmarks.
+In Figures 5 and 6 we report pass@10 scores for all models on HumanEval and HumanEval+ benchmarks.
 Figure 5: HumanEval pass@10 scores for quantized Llama-3.1-Instruct models.
@@
 A.4 Academic Benchmarks
-In Tables 10 and 11 we report accuracy recoveries per-task across academic benchmarks.
+In Tables 9 and 10 we report accuracy recoveries per-task across academic benchmarks.
 Table 9: Accuracy recoveries in percentages (%) for each task in the Open LLM Leaderboard V1 benchmark.

### v3

@@
 Accuracy-Performance Trade-Offs in LLM Quantization
-Eldar Kurtic1,2, Alexandre Marques1, Shubhra Pandit1, Mark Kurtz1, Dan Alistarh1,2 1Red Hat AI, 2Institute of Science and Technology Austria Correspondence: ekurtic@redhat.com, dalistar@redhat.com
+Eldar Kurtić1,2, Alexandre Marques1, Shubhra Pandit1, Mark Kurtz1, Dan Alistarh1,2 1Red Hat AI, 2Institute of Science and Technology Austria Correspondence: ekurtic@redhat.com, dalistar@redhat.com
 1 Introduction
@@
 Figure 1: Text similarity metrics comparing the outputs of quantized Llama-3.1-Instruct models to full-precision baselines. We refer to W8A8-FP as FP8, W8A8-INT as INT8, and W4A16-INT as INT4.
-4.2 Accuracy on Real-World Benchmarks
+4.2 Real-World Benchmarks
 While academic benchmarks offer structured evaluations, real-world benchmarks better capture model performance in dynamic environments.
@@
 See Appendix A.1 for additional results.
+4.3 Reasoning Benchmarks
+Given the recent rise in popularity of reasoning abilities of LLMs, we also focus on the popular DeepSeek-R1-Distill DeepSeek-AI (2025) models.
+These models have been fine-tuned through the process of distillation for improved reasoning capabilities.
+To assess their reasoning performance, we focus on the challenging and widely recognized reasoning benchmarksi through LightEval Habib et al.
+(2023): AIME 2024, MATH-500 Lightman et al.
+(2023), and GPQA-Diamond Rein et al.
+(2024).
+Following DeepSeek’s recommendations for text generation, we use sampling with a temperature of 0.6 and top-p of 0.95, generating 20 responses per query to estimate the pass@1 score.
+The repetitive sampling was important to estimate an accurate average performance for the benchmarks due to high variance across the relatively small datasets.
+As can be seen from the results in Table 4, the conclusions from the previous sections with academic and real-world benchmarks still hold: when quantization is properly tuned and configured, quantized models perform very competitively with their unquantized (BF16) baselines, recovering on average >99% accuracy except for the smallest models at INT4 which exhibit a bit larger but reasonable drops.
+Table 4: Detailed per-task and per-model breakdown of accuracy on the popular reasoning benchmarks across all quantized variants of DeepSeek-R1-Distill models from both Llama and Qwen families.
+DeepSeek-R1-Distill Recovery % Average Score AIME24 pass@1 MATH-500 pass@1 GPQA-Diamond pass@1
+Llama-8B BF16 100.0 62.9 49.3 ± 6.4 90.2 ± 1.2 49.3 ± 3.1
+W8A8-FP 100.6 63.3 50.8 ± 9.0 90.2 ± 1.1 48.7 ± 2.5
+W8A8-INT 99.6 62.7 49.1 ± 6.2 90.0 ± 1.0 48.9 ± 2.0
+W4A16-INT 97.2 61.1 46.3 ± 6.9 89.9 ± 1.1 47.1 ± 2.6
+Llama-70B BF16 100.0 76.2 67.8 ± 7.2 95.3 ± 0.7 65.6 ± 2.3
+W8A8-FP 100.3 76.5 69.2 ± 6.5 95.1 ± 0.5 65.2 ± 2.4
+W8A8-INT 99.7 76.0 67.8 ± 6.4 95.3 ± 0.5 65.0 ± 1.8
+W4A16-INT 98.3 75.0 65.6 ± 5.3 95.2 ± 0.6 64.0 ± 2.8
+Qwen-32B BF16 100.0 76.3 69.8 ± 4.9 95.1 ± 0.6 64.1 ± 2.1
+W8A8-FP 99.0 75.6 68.5 ± 4.0 95.3 ± 0.7 62.9 ± 2.6
+W8A8-INT 99.6 76.0 68.2 ± 5.1 95.0 ± 0.8 64.8 ± 2.6
+W4A16-INT 99.5 75.9 68.8 ± 4.2 95.0 ± 0.5 63.8 ± 1.7
+Qwen-14B BF16 100.0 73.6 66.7 ± 5.1 94.7 ± 0.7 59.4 ± 2.3
+W8A8-FP 101.0 74.3 68.1 ± 5.8 94.6 ± 0.6 60.1 ± 2.9
+W8A8-INT 99.4 73.1 66.3 ± 7.1 94.7 ± 0.7 58.3 ± 2.0
+W4A16-INT 99.0 72.8 66.0 ± 6.3 95.0 ± 0.5 57.5 ± 2.1
+Qwen-7B BF16 100.0 65.8 53.2 ± 6.4 93.7 ± 0.8 50.5 ± 2.8
+W8A8-FP 99.9 65.7 53.2 ± 7.5 93.6 ± 0.7 50.3 ± 2.0
+W8A8-INT 100.7 66.3 55.2 ± 4.9 93.0 ± 1.1 50.7 ± 3.5
+W4A16-INT 98.3 64.7 50.9 ± 7.8 93.3 ± 1.1 49.8 ± 2.8
+Qwen-1.5B BF16 100.0 50.0 30.1 ± 5.3 84.7 ± 1.1 35.4 ± 3.0
+W8A8-FP 100.3 50.2 29.8 ± 5.6 84.7 ± 1.3 35.9 ± 3.3
+W8A8-INT 96.9 48.5 26.7 ± 6.3 84.4 ± 1.1 34.4 ± 2.8
+W4A16-INT 93.5 46.8 24.6 ± 5.1 82.5 ± 1.1 33.2 ± 3.4
 4.4 Text Similarity Investigation
@@
 A.4 Academic Benchmarks
-In Tables 9 and 10 we report accuracy recoveries per-task across academic benchmarks.
+In Tables 10 and 11 we report accuracy recoveries per-task across academic benchmarks.
 Table 10: Accuracy recoveries in percentages (%) for each task in the Open LLM Leaderboard V1 benchmark.